### v1

# Ramsey Numbers in Kneser Graphs

Abstract
We define the $r$-Kneser Ramsey number $R^{\textrm{KG}}_{r}(s,t)$ as the minimum integer $n$ such that every red/blue edge-coloring of the Kneser graph $\textrm{KG}(n,r)$ contains a red $s$-clique or a blue $t$-clique. We obtain general bounds on the numbers $R^{\textrm{KG}}_{r}(s,t)$, and make progress on two related Ramsey-type problems, one raised by Holmsen, Hrusak, and Roldán-Pensado, and the other posted by Pálvölgyi.
1 Introduction
The Ramsey number $R(s,t)$ is the minimum integer $n$ such that any red/blue edge-coloring of the complete graph $K_{n}$ contains either a red $K_{s}$ or a blue $K_{t}$. Ramsey numbers and their many different variations have been extensively studied, and they form one of the main pillars of extremal combinatorics. In this paper we consider a variation on Ramsey numbers in which the host graph is a Kneser graph rather than the complete graph. The Kneser graph $\textrm{KG}(n,r)$ is the graph whose vertex set is the set $n]\choose r$} of all $r$-subsets of $[n]$, and two $r$-sets form an edge if they are disjoint. We define the Kneser Ramsey numbers as follows.
Definition 1.
Let $r\geq 1$ be an integer. The $r$-Kneser Ramsey number $R^{\textrm{KG}}_{r}(s,t)$ is the minimum integer $n$ such that every red/blue edge-coloring of $\textrm{KG}(n,r)$ contains a red $K_{s}$ or a blue $K_{t}$.
Note that $R^{\textrm{KG}}_{1}(s,t)=R(s,t)$, so this definition generalizes the notion of Ramsey numbers.
Our motivation to investigate Kneser Ramsey numbers originated with the following fascinating Ramsey-type question, which was asked by Michael Hrusak for $\alpha=1/2$ and then by Andreas Holmsen and Edgardo Roldán-Pensado for all $\alpha$ [4]:
Problem 2 (Holmsen - Hrusak - Roldán-Pensado).
Fix $\alpha\in(0,1)$ and an integer $t>1$. Does there exist a large enough $n$ (depending on $\alpha$ and $t$) such that in every red/blue edge-coloring of the complete graph on vertex set $V={[n]\choose\alpha n}$ there is a monochromatic clique on vertex set $U\subset V$, with the property that the vertex-covering number of the hypergraph $([n],U)$ is at least $t$?
Although Problem 2 is not about Kneser graphs, one observes that a positive answer will follow if in every red/blue edge-coloring of $\textrm{KG}(n,\alpha n)$ there is a monochromatic clique of size $t$. That is, if $R^{\textrm{KG}}_{\alpha n}(t,t)\leq n$ for some $n$, then we get a positive answer to Problem 2.
However, Problem 2 is interesting even when $\textrm{KG}(n,\alpha n)$ is empty. In fact, Holmsen and Roldán-Pensado were most interested in the special case when $\alpha=k/(2k-1)$ and $t=2$, as formulated below.
Problem 3.
Let $V={[2k-1]\choose k}$. Is it true that for some large enough $k$, in every red/blue edge-coloring of the complete graph on vertex set $V$ there is a monochromatic triangle $ABC$ with $A,B,C\in V$ and $A\cap B\cap C=\emptyset$?
Perhaps a more approachable problem is the following.
Problem 4.
Determine the smallest number $\beta>1$ such that for some large enough $k$, in every red/blue edge-coloring of the complete graph on vertex set $V={[\beta k]\choose k}$ there is a monochromatic triangle $ABC$ with $A,B,C\in V$ and $A\cap B\cap C=\emptyset$.
It is easy to see that $\beta\leq 3$. Indeed, let $S_{1},\dots,S_{6}\in{[3k]\choose k}$ be six sets with the property that for every $\in[3k], tt$ belongs to exactly two of the sets $S_{1},\dots,S_{6}$. (It is easy to construct such sets $S_{1},\dots,S_{6}$: take the six distinct sets in two disjoint perfect matchings in the $k$-uniform hypergraph on vertex set $[3k]$.) Now consider a red/blue edge-coloring of the complete graph on vertices $S_{1},\dots,S_{6}$. Since $R(3,3)=6$, this graph contains a monochromatic triangle $S_{i}S_{j}S_{k}$, and $S_{i}\cap S_{j}\cap S_{k}=\emptyset$ for every distinct $i,j,k\in[6]$.
In this paper we make progress on Problems 2 and 4 in two different ways. First, in Section 2 we improve the upper bound on $\beta$ to $7/3$.
Theorem 5.
Let $k\geq 12$ be an integer divisible by 6. Then in every red/blue edge-coloring of the complete graph on vertex set $V={[\frac{7}{3}k]\choose k}$ there is a monochromatic triangle $ABC$ with $A,B,C\in V$ and $A\cap B\cap C=\emptyset$.
Second, in the same section, we give a positive answer to Problem 2 when $\alpha<1/4$ and $t=3$.
Theorem 6.
For any $0<\alpha<\frac{1}{4}$, there exists some large enough $n$ such that any red/blue edge-coloring of $\textrm{KG}(n,\alpha n)$ contains a monochromatic triangle.
Note that for $\alpha=\frac{1}{4}$, for $n=8$ there exists a red/blue edge-coloring of $\textrm{KG}(n,\alpha n)$ with no monochromatic triangle, as implied by Theorem 12.
We then continue with a systematic study of Kneser Ramsey numbers for various parameters $r,s,t$. The following trivial upper bound on the $r$-Kneser Ramsey number follows from Ramsey’s theorem:
$R^{\textrm{KG}}_{r}(s,t)\leq R(s,t)\cdot r.$
Indeed, when $n\geq R(s,t)\cdot r$, the graph $\textrm{KG}(n,r)$ contains a clique of size $R(s,t)$, and any red/blue edge-coloring of it contains a red $K_{s}$ or a blue $K_{t}$. This in particular shows the existence of the number $R^{\textrm{KG}}_{r}(s,t)$. Note also that if $n<kr$, then $\textrm{KG}(n,r)$ does not contain a clique of size $k$, and thus we have a trivial lower bound
$R^{\textrm{KG}}_{r}(s,t)\geq r\cdot\max\{s,t\}.$
In order to prove Theorem 6, we prove an auxiliary result (Theorem 17) that in particular implies the following bound on $R^{\textrm{KG}}_{r}(s,t)$.
Theorem 7.
Fix $s,t\geq 3$ and suppose $R(s-1,t)\geq R(s,t-1)$. Set M=R(s1,t)M=R(s-1,t) and m=R(s,t1)m=R(s,t-1). We have
$R^{\textrm{KG}}_{r}(s,t)\leq(M+1)r+m-1.$
In particular,
$R^{\textrm{KG}}_{r}(s,s)\leq(R(s,s-1)+1)r+R(s,s-1)-1.$
In Section 3 we establish improved upper bounds on $R^{\textrm{KG}}_{r}(s,t)$ for certain families of $s,t$.
Theorem 8.
For $s\geq 4$, we have
$R^{\textrm{KG}}_{r}(s-1,s)\leq rR(s-2,s)+2(R(s-2,s-1)-1)+2r.$
Theorem 9.
For $t\geq 5$, we have $R^{\textrm{KG}}_{2}(3,t)\leq 2(R(3,t-2))+2t+1.$
We then derive a general lower bound, and an improved lower bound for $s=t=3$ and $r\geq 3$.
Proposition 10.
For $r\geq 1$ and $s,t\geq 2$, we have $R^{\textrm{KG}}_{r}(s,t)\geq R(s,t)+2r-2.$
Theorem 11.
For $r\geq 2$, we have $R^{\textrm{KG}}_{r}(3,3)\geq 3r+2.$
Using a computer code, we also obtain an exact value on $R^{\textrm{KG}}_{2}(3,3)$. The computational proof of the next theorem is discussed in Subsection 3.3.
Theorem 12.
We have $R^{\textrm{KG}}_{2}(3,3)=9$ and $R^{\textrm{KG}}_{3}(3,3)\leq 13$.
Some of our results are summarized in Table 1 (for $r=2$) and Table 2 (for $r=3$). In each cell of the tables, the first number listed is the lower bound and the second is the upper bound. The subscripts on the numbers are the theorems in which we obtain the bound. For example, we have $11\leq R^{\textrm{KG}}_{2}(3,4)\leq 16$, where the lower bound follows from Theorem 10 and the upper bound is proven in Theorem 8. If there is only one number, then it is an exact value. Note that $R^{\textrm{KG}}_{r}(s,t)=R^{\textrm{KG}}_{r}(t,s)$.
Table 1: Bounds on $R^{\textrm{KG}}_{2}(s,t)$.
$s$ $t$ $t=3$ $t=4$ $t=5$ $t=6$
$s=3$ 912 1110, 168 1610, 239 2010, 319
$s=4$ 2010, 287 2710, 488 3810, 697
$s=5$ 4510, 767 6010, 1348
$s=6$ 10410, 2627
Table 2: Bounds on $R^{\textrm{KG}}_{3}(s,t)$.
$s$ $t$ $t=3$ $t=4$ $t=5$ $t=6$
$s=3$ 1111, 1312 1310, 247 1810, 347 2210, 507
$s=4$ 2210, 387 2910, 707 4010, 957
$s=5$ 4710, 1027 6210, 1877
$s=6$ 10610, 3507
We conclude this paper with some progress on a Ramsey-type question that was posted by Dömötör Pálvölgyi on mathoverflow in 2016 [6]. Here an induced $\textrm{KG}(3r,r)$ is a subgraph of $\textrm{KG}(n,r)$ induced on vertex set $S\choose r$}, where $S\subset[n]$ is a set of size $3r$.
Problem 13 (Pálvölgyi [6]).
For $r>1$, determine the smallest $N\geq 3r$ such that in any red/blue edge-coloring of $\textrm{KG}(N,r)$ there exists either a red induced $\textrm{KG}(3r,r)$ or a blue triangle.
In Section 4 we prove a lower bound, depending on Ramsey numbers, for the number $N$ in Pálvölgyi’s problem. For integers $r,s,t$, let $R_{r}(s,t)$ be the minimum integer $n$ such that in any red/blue coloring of the edges of the complete $r$-uniform hypergraph on $n$ vertices $K^{r}_{n}$, there is a red $s$-clique or a blue $t$-clique.
Theorem 14.
We have $N\geq R_{r}(3r,3r)$. That is, for $n=R_{r}(3r,3r)-1$, there is a red/blue coloring of $\textrm{KG}(n,r)$ containing no red induced $\textrm{KG}(3r,r)$ and no blue triangle.
For example, for $r=2$ we get $N\geq 102$, as $R_{2}(6,6)=R(6,6)\geq 102$ [3].
Many of our proofs use a theorem of Alon, Frankl, and Lovász [1].
Theorem 15 (Alon - Frankl - Lovász [1]).
Let $k_{1}\geq k_{2}\geq\dots\geq k_{t}\geq 2$ and suppose $n\geq k_{1}r+\sum_{i=2}^{t}(k_{i}-1)$. Let $F_{1},F_{2},\dots,F_{t}$ be a partition of $\binom{[n]}{r}$. Then, for some $i\in[t]$, $F_{i}$ contains $k_{i}$ pairwise disjoint sets.
Observe that when $k_{1}=k_{2}=\cdots=k_{t}=2$, Theorem 15 recovers the chromatic number of the Kneser graph. When $t=2$, we apply a simpler form.
Theorem 16 (Alon - Frankl - Lovász [1]).
Let $n\geq(r+1)k-1$ and let $F_{1},F_{2}$ be a partition of $\binom{[n]}{r}$. Then, either $F_{1}$ or $F_{2}$ contains $k$ pairwise disjoint sets.
2 The Holmsen - Hrusak - Roldán-Pensado problem
We start with the proof of Theorem 5, showing that $7/3$ is an upper bound on $\beta$ in Problem 4.
Proof of Theorem 5. Let $V=[3k/2]\cup T$, where $T$ is a set of $5k/6$ vertices disjoint from $[3k/2]$. Then $|V|=7k/3$. Suppose we have a red/blue edge-coloring of the complete graph $G$ on $V\choose k$}.
Consider the triangle $S_{1}S_{2}S_{3}$ in $G$, where $S_{1}=\{1,\dots,k\}$, $S_{2}=\{1,\dots,k/2\}\cup\{k+1,\dots,3k/2\}$, and $S_{3}=\{k/2+1,\dots,3k/2\}$. Since $S_{1}\cap S_{2}\cap S_{3}=\emptyset$, if $S_{1}S_{2}S_{3}$ is monochromatic we are done. Otherwise, it must have a red edge and a blue edge. Without loss of generality, $S_{1}S_{2}$ is red, and $S_{1}S_{3}$ is blue. Note that $S_{1}\cap S_{i}\subset[k]$ for all $i$.
Let $T_{1},T_{2},T_{3}$ be pairwise disjoint subsets of $\{k+1,\dots,3k/2\}$ of size $k/6$ each. Let $F_{i}=T\cup T_{i}$ for $i\in[3]$.
Assume first that the triangle $F_{1}F_{2}F_{3}$ is monochromatic. If it is monochromatic blue, then consider the induced graph of $G$ on vertex set $\{S_{1},S_{2},F_{1},F_{2},F_{3}\}$. Observe that if either $S_{1}$ or $S_{2}$ has at least two blue edges to vertices in $\{F_{1},F_{2},F_{3}\}$, then we get a blue triangle of the form $S_{i}F_{j}F_{l}$, and since $S_{i}\cap F_{j}\cap F_{l}=S_{i}\cap T=\emptyset$, we are done. Otherwise, both $S_{1}$ and $S_{2}$ have at least two red edges to vertices in $\{F_{1},F_{2},F_{3}\}$, so by the pigeonhole principle we get a red triangle of the form $S_{1}S_{2}F_{i}$. But now $(S_{1}\cap S_{2}\cap F_{i})\subset([k]\cap F_{i})=\emptyset$, and we are done. If the triangle $F_{1}F_{2}F_{3}$ is monochromatic red, the argument is symmetric when we consider instead the induced graph on vertex set $\{S_{1},S_{3},F_{1},F_{2},F_{3}\}$.
So suppose the triangle $F_{1}F_{2}F_{3}$ is not monochromatic. Then without loss of generality $F_{1}F_{2}$ is a red edge and $F_{1}F_{3}$ is a blue edge.
Consider the induced graph on vertex set $\{H_{1},\dots,H_{6}\}$, where $H_{i}=[k-1]\cup\{k+i-1\}$ for $i\in[6]$. Since $R(3,3)=6$, this graph has a monochromatic triangle, say $H_{1}H_{2}H_{3}$. Note that $H_{i}\cap H_{j}=[k-1]$ for $i\neq j$. Moreover, since $k/2\geq 6$ we have $H_{i}\cap T=\emptyset$ for all $i\in[6]$.
If the triangle $H_{1}H_{2}H_{3}$ is monochromatic blue, then consider the induced graph of $G$ on vertex set $\{F_{1},F_{2},H_{1},H_{2},H_{3}\}$. Observe that if $F_{1}$ or $F_{2}$ has at least two blue edges to vertices in $\{H_{1},H_{2},H_{3}\}$, then we get a blue triangle of the form $F_{i}H_{j}H_{l}$, and since $F_{i}\cap H_{j}\cap H_{l}=F_{i}\cap[k-1]=\emptyset$, we are done. Otherwise, both $F_{1}$ and $F_{2}$ have at least two red edges to vertices in $\{H_{1},H_{2},H_{3}\}$, so by the pigeonhole principle we get a red triangle of the form $F_{1}F_{2}H_{i}$. But now $(F_{1}\cap F_{2}\cap H_{i})\subset(T\cap H_{i})=\emptyset$, and again we are done.
Finally, if the triangle $H_{1}H_{2}H_{3}$ is monochromatic red, we have a symmetric argument by considering instead the induced graph on vertex set $\{F_{1},F_{3},H_{1},H_{2},H_{3}\}$. ∎
Our next goal is to prove Theorem 6, which gives a positive answer to Problem 2 in the case $\alpha<1/4$ and $t=3$. In order to prove the theorem, we need the following auxiliary result.
Theorem 17.
Fix $s,t\geq 3$ and suppose $R(s-1,t)\geq R(s,t-1)$. Set M=R(s1,t)M=R(s-1,t) and m=R(s,t1)m=R(s,t-1). Let $0<\varepsilon<\frac{1}{M+1}$ and set $\alpha=\frac{1}{M+1}-\varepsilon$. Then, there exists some $n\in\mathbb{N}$ such that $\alpha n\in\mathbb{N}$ and $\textrm{KG}(n,\alpha n)$ contains a red $K_{s}$ or a blue $K_{t}$.
Proof of Theorem 17.
Fix $r\geq 2$ and set Nr=Mr+m1N-r=Mr+m-1. First, we will show that in any red/blue coloring of $G=\textrm{KG}(N,r)$, there is a red $K_{s}$ or a blue $K_{t}$. Suppose for contradiction that there is some coloring $c:E(G)\to\{R,B\}$ of $G$ without a red $K_{s}$ or blue $K_{t}$. Let $v=[N]\setminus[N-r]\in V(G)$. Then, $\binom{[N-r]}{r}=\binom{[Mr+m-1]}{r}$ can be partitioned into $N_{R}(v)$ and $N_{B}(v)$. Since $c$ contains no red $K_{s}$ and no blue $K_{t}$, then $G[N_{R}(v)]$ contains no red $K_{s-1}$ and no blue $K_{t}$. Similarly, $G[N_{B}(v)]$ contains no red $K_{s}$ and no blue $K_{t-1}$. In particular, $G[N_{R}(v)]$ contains no $K_{M}$ and $G[N_{B}(v)]$ contains no $K_{m}$. However, by Theorem 16, setting $F_{1}=N_{R}(v)$ and $F_{2}=N_{B}(v)$, either $N_{R}(v)$ contains a clique of size $M$ or $N_{B}(v)$ contains a clique of size $m$, a contradiction.
For the general case, fix $0<\varepsilon<\frac{1}{M+1}$ and set $\alpha=\frac{1}{M+1}-\varepsilon$. Choose $r\in\mathbb{Z}$ such that $\frac{r}{(M+1)r+m-1}\geq\alpha$. Set Nr=Mr+m1N-r=Mr+m-1 and $n=\frac{r}{\alpha}$. Then, by choice of $r$, observe that $N\leq n$. Therefore, since $\textrm{KG}(N,\alpha n)\subseteq\textrm{KG}(n,\alpha n)$, any red-blue coloring of $\textrm{KG}(n,\alpha n)$ contains a red $K_{s}$ or a blue $K_{t}$. ∎
Note that the general bound on $R^{\textrm{KG}}_{r}(s,t)$ in Theorem 7 is proven in the first paragraph of the proof of Theorem 17 above.
We are now ready to prove Theorem 6.
Proof of Theorem 6.
We apply Theorem 17 with $s=t=3$. This gives $M=m=R(2,3)=3$. For any $0<\varepsilon<\frac{1}{4}$, we have $\alpha=\frac{1}{4}-\varepsilon$, and by Theorem 17 there exists some $n\in\mathbb{N}$ with $\alpha n\in\mathbb{N}$ such that any red/blue coloring of the edges of $\textrm{KG}(n,\alpha n)$ contains a monochromatic triangle. ∎
3 General bounds on Kneser Ramsey numbers
In this section we prove Theorems 8, 9, 11, and 12, and Proposition 10.
3.1 Upper bounds
Proof of Theorem 8.
We would like to show that for $s\geq 4$,
$R^{\textrm{KG}}_{r}(s-1,s)\leq rR(s-2,s)+2(R(s-2,s-1)-1)+2r.$
Let N=rR(s2,s)+2R(s2,s1)+2rN=rR(s-2,s)+2R(s-2,s-1)+2r, $G=\textrm{KG}(N,r)$ and let $c:E(G)\to\{R,B\}$ be a red/blue edge-coloring of $G$. For sake of contradiction, suppose there is no red $K_{s-1}$ nor blue $K_{s}$. If there is no blue edge in $G$, then we are done. So, set $u=\{N,N-1,\dots,N-r+1\}\in V(G)$ and $v=\{N-r,N-r-1,\dots,N-2r+1\}\in V(G)$. Without loss of generality, suppose $c(u,v)=B$. Then, $N(u)\cap N(v)\cong G^{\prime}=\textrm{KG}(N-2r,r)$. We will partition the vertices of $G^{\prime}$ into 4 sets. For $(X,Y)\in\{(R,R),(R,B),(B,R),(B,B)\}$, we denote by $N_{XY}$ the set of vertices $w$ in $G^{\prime}$ such that $c(w,u)=X$ and $c(w,v)=Y$.
Observe that $N_{BR}$ must contain no red $K_{s-2}$ or we may find a red $K_{s-1}$ along with $v$, and $N_{BR}$ must contain no blue $K_{s-1}$ or we may find a blue $K_{s}$ along with $u$. Thus $\omega(G^{\prime}[N_{BR}])\leq R(s-2,s-1)-1$. Similarly, $N_{RR}\cup N_{RB}$ must contain no red $K_{s-2}$ or blue $K_{s}$, so $\omega(G^{\prime}[N_{RR}\cup N_{RB}])\leq R(s-2,s)-1$. Furthermore, $N_{BB}$ must contain no red $K_{s-1}$ and no blue $K_{s-2}$, as otherwise we can find a blue $K_{s}$ along with the edge $uv$. Hence $\omega(G^{\prime}[N_{BB}])\leq R(s-2,s-1)-1$.
However, invoking Theorem 15 with $F_{1}=N_{RR}\cup N_{RB}$ ($k_{1}=R(s-2,s)$), $F_{2}=N_{BR}$ ($k_{2}=R(s-2,s-1)$), and $F_{3}=N_{BB}$ ($k_{3}=R(s-2,s-1)$), since $k_{1}\cdot r+(k_{2}-1)+(k_{3}-1)=N-2r$, either $F_{1}$ contains a clique on R(s2,s)R(s-2,s) vertices, $F_{2}$ contains a clique on R(s2,s1)R(s-2,s-1) vertices, or $F_{3}$ contains a clique on R(s2,s1)R(s-2,s-1) vertices. In any case, we are guaranteed either a red $K_{s-1}$ or a blue $K_{s}$, a contradiction. Therefore, $R^{\textrm{KG}}_{r}(s-1,s)\leq N$. ∎
Theorem 9 gives an improvement when $r=2$ and $s=3$.
Proof of Theorem 9.
For sake of contradiction, fix N>2(R(3,t2))+2tN>2(R(3,t-2))+2t, let $G=\textrm{KG}(N,2)$, $c:E(G)\to\{R,B\}$ be a red/blue edge-coloring of $G$, and suppose there is no red $K_{3}$ or blue $K_{t}$. Note that if there are no blue edges, then there must be a red $K_{3}$, and we have arrived at our contradiction.
Hence, without loss of generality, we may assume $c(\{1,2\},\{3,4\})=B$. Next, we look at $N(\{1,2\})\cap N(\{3,4\})\cong G^{\prime}=\textrm{KG}(N-4,2)$, and we will partition the vertices of $G^{\prime}$ into $4$ sets. For $(X,Y)\in\{(R,R),(R,B),(B,R),(B,B)\}$, we denote by $N_{XY}$ the set of vertices $v$ in $G^{\prime}$ such that $c(v,\{1,2\})=X$ and $c(v,\{3,4\})=Y$.
First, observe that for any two vertices $u,v\in N_{RB}$ with $uv\in E(G^{\prime})$, $c(u,v)=B$ (otherwise there is a red $K_{3}$ with $u,v,\{1,2\}$, a contradiction). Similarly, for each case of $u,v\in N_{RR}$, $u,v\in N_{BR}$, and $u\in N_{RR},v\in N_{RB}$, we have $c(u,v)=B$. These observations allow us to bound the clique numbers of $N_{RR}\cup N_{RB}$ and $N_{BR}$. In particular, $\omega(G^{\prime}[N_{RR}\cup N_{RB}])\leq t-1$ since all edges are blue and we cannot have a blue $K_{t}$. Additionally, $\omega(G^{\prime}[N_{BR}])\leq t-2$ since all edges are blue, and if there were a blue $K_{t-1}$, together with vertex $\{1,2\}$, we would have a blue $K_{t}$.
Finally, $\omega(G^{\prime}[N_{BB}])<R(t-2,3)$ since otherwise, there would be either a copy of a red $K_{3}$ (a contradiction), or a blue copy of $K_{t-2}$, which together with the blue edge $(\{1,2\},\{3,4\})$ would be a blue $K_{t}$, another contradiction.
By Theorem 15, with $F_{1}=N_{RR}\cup N_{RB}$ ($k_{1}=R(t-2,3))$, $F_{2}=N_{BR}$ ($k_{2}=t$), and $F_{3}=N_{BB}$ ($k_{3}=t-1$), since $N-4\geq 2(R(t-2,3))+(t-1)+(t-1-1)$, either $F_{1}$ contains a $K_{R(t-2,3)}$, $F_{2}$ contains a $K_{t}$, or $F_{3}$ contains a $K_{t-1}$, a contradiction in all cases. This final contradiction proves the result. ∎
3.2 Lower bounds
The lower bound in Proposition 10 takes advantage of the chromatic number of Kneser graphs. Lovász  [5] proved that $\chi(\textrm{KG}(n,r))=n-2r+2$, which was one of the earliest applications of topological methods to combinatorics.
Proof of Proposition 10.
Let $n=R^{\textrm{KG}}_{r}(s,t)$ and set $G=\textrm{KG}(n,r)$. Suppose $\chi(G)=n-2r+2<R(s,t)$ and let $c:V(G)\to[n-2r+2]$ be a proper vertex-coloring of $G$. Consider the complete graph $G^{\prime}$ on n2r+2n-2r+2 vertices with $V(G^{\prime})=[n-2r+2]$ and let $C^{\prime}:E(G^{\prime})\to\{R,B\}$ be a red/blue edge-coloring of $G^{\prime}$ avoiding a red copy of $K_{s}$ and a blue copy of $K_{t}$.
Now, consider the following red/blue edge-coloring of $G$: For $uv\in E(G)$, color $uv$ with $C^{\prime}(c(u),c(v))$. We claim that this coloring has no red copy of $K_{s}$ or blue copy of $K_{t}$. Suppose there is a red copy of $K_{s}$ with vertices $\{v_{1},v_{2},\dots,v_{s}\}$. Observe that for all $i\neq j$, $c(v_{i})\neq c(v_{j})$ since $c$ is a proper vertex-coloring and $v_{i}v_{j}\in E(G)$. But now, the color of $v_{i}v_{j}$ is $C^{\prime}(c(v_{i}),c(v_{j}))=R$. This implies that the graph induced by the vertices $\{c(v_{1}),c(v_{2}),\dots,c(v_{s})\}$ is a red copy of $K_{s}$, a contradiction to our choice of $C^{\prime}$. Therefore, $n-2r+2\geq R(s,t)$. ∎
Proof of Theorem 11.
We wish to show there exists a red/blue coloring $c$ of $G=\textrm{KG}(3r+1,r)$ with no monochromatic triangle. To this end, we partition $V(G)$ as follows:
$X$ $=\{x\in V(G):3r+1\in x\},$
$Y$ $=\{x\in V(G)\setminus X:3r\in x\},$
$Z$ $=V(G)-Y-X\simeq V(\textrm{KG}(3r-1,r)).$
Observe that the induced subgraph $G[Z]$ of $G$ on $Z$ is triangle-free. Also, $X$ and $Y$ are independent sets with no triangles contained in $G[X\cup Y]$. Therefore, the only triangles will have either two vertices from $Z$ and one from $X\cup Y$ or one vertex from each of $X,Y,Z$. Setting $c(E(G[Z]))=B$ and $c(E(X,Y))=B$ with all other edges colored $R$ (see Figure 1), we obtain a red/blue coloring with no monochromatic triangles. ∎
XXYYZZ
Figure 1: The coloring in 11
3.3 Computationally-obtained bounds
To obtain the upper and lower bounds in Theorem 12 we wrote a program in Python, which for (small enough) specified parameters $n,r,s,t$, runs over all possible red/blue edge-colorings of $\textrm{KG}(n,r)$ and either provides as output a coloring that contains no red $K_{s}$ and no blue $K_{t}$ (which we call “good”), or declares that a good coloring does not exists.
We do this by reformulating the question as a satisfiability problem. For example, for $R^{\textrm{KG}}_{r}(3,3)$, we look for a red/blue edge-coloring $c$ of $G=\textrm{KG}(n,r)$ with no monochromatic triangle. Suppose $E(G)=\{e_{1},\dots,e_{|E(G)|}\}$. For $1\leq i\leq|E(G)|$, we may assign a boolean variable $v_{i}$ corresponding to $e_{i}$, where
$v_{i}=\begin{cases}1\text{ (True)}&\text{if }c(e_{i})=R,\\ 0\text{ (False)}&\text{if }c(e_{i})=B.\end{cases}$
Then, for a triangle $T=\{e_{1},e_{2},e_{3}\}$ of $G$, $T$ is monochromatic if and only if the following boolean clause is true:
$(v_{1}\land v_{2}\land v_{3})\lor(\neg v_{1}\land\neg v_{2}\land\neg v_{3}).$
Negating this, we get that $T$ is not monochromatic if and only if the following boolean clause, $B_{T}$, is true:
$B_{T}=(v_{1}\lor v_{2}\lor v_{3})\land(\neg v_{1}\lor\neg v_{2}\lor\neg v_{3}).$
Let $\{T_{1},\dots,T_{t}\}$ be the set of all triangles in $G$. Then, there is a good edge-coloring of $G$ if and only if there exists an assignment $(v_{1},\dots,v_{|E(G)|})\in\{0,1\}^{|E(G)|}$ such that the clause $B=\bigwedge_{i=1}^{t}B_{T_{i}}$ is satisfiable (True).
Our program builds the clause $B$ and then checks its satisfiability.
The proof of Theorem 12 follows because our program found a good coloring of $\textrm{KG}(8,2)$ (this coloring is given explicitly in Appendix A), but declared that the graphs $\textrm{KG}(9,2)$ and $\textrm{KG}(13,3)$ do not have good colorings.
Our code is given in Appendix B.
4 Pálvölgyi’s problem
We begin with a general discussion on monochromatic induced Kneser subgraphs, which will be useful in making progress on Problem 13.
Let $r>1$ be an integer and let $\textrm{KG}(s,r)$ and $\textrm{KG}(t,r)$ be Kneser graphs. Define the induced $r$-Kneser Ramsey number $R^{\textrm{KG}}_{r,ind}(s,t)$ to be the minimum $n$ (if it exists) such that every red/blue edge-coloring of $\textrm{KG}(n,r)$ contains an induced red $\textrm{KG}(s,r)$ or an induced blue $\textrm{KG}(t,r)$.
Theorem 18.
For $r\geq 2$, $R^{\textrm{KG}}_{r,ind}(s,t)$ does not exists; that is, there exists a red/blue edge-coloring of $\textrm{KG}(n,r)$ with no red induced $\textrm{KG}(s,r)$ or blue induced $\textrm{KG}(t,r)$ .
In the proof of Theorem 18 we use the set-system theorem of Bollobás.
Theorem 19 (Bollobás [2]).
Let $(A_{1},B_{1}),\dots,(A_{m},B_{m})$ be pairs of sets with $|A_{i}|=a$ and $|B_{i}|=b$ for every $i$. Suppose that $A_{i}\cap B_{i}=\emptyset$ for $1\leq i\leq m$, and $A_{i}\cap B_{j}\neq\emptyset$ for $i\neq j$. Then $m\leq\binom{a+b}{a}$. Furthermore, if $m=\binom{a+b}{a}$ then there is some set $S$ of cardinality $a+b$ such that the sets $A_{i}$ are all the subsets of $S$ of size $a$, and $B_{i}=S\setminus A_{i}$ for all $i$.
Proof of Theorem 18.
We first show $R^{\textrm{KG}}_{r,ind}(2r,2r)$ does not exist. Observe that $\textrm{KG}(2r,r)$ is a matching of size $\frac{1}{2}\binom{2r}{r}$.
Now, in any $\textrm{KG}(n,r)$ with $n\geq 2r$, we claim that every maximum induced matching $M$ has vertex set $\binom{S}{r}$, where $|S|=2r$. To this end, let $M$ be an induced matching with edge set $E=\{(X_{i},Y_{i})\colon 1\leq i\leq m\}$. Note that we must have $X_{i}\cap Y_{i}=\emptyset$ for all $i$, and $X_{i}\cap Y_{j}\neq\emptyset$ for all $i\neq j$ in order for $M$ to be an induced matching. Let $A_{i}=X_{i}$ for $1\leq i\leq m$, $A_{i}=Y_{i-m}$ for $m+1\leq i\leq 2m$, and let $B_{i}=Y_{i}$ for $1\leq i\leq m$, $B_{i}=X_{i-m}$ for $m+1\leq i\leq 2m$. Then the set pair system $(A_{i},B_{i})$, $1\leq i\leq 2m$ satisfies $A_{i}\cap B_{i}=\emptyset$ for all $i$ and $A_{i}\cap B_{j}\neq\emptyset$ for $i\neq j$.
Thus, we may invoke Theorem 19, with $a=b=r$, and
$2\cdot|E(M)|=2m\leq\binom{2r}{r}$.
In fact, since we know there exists an induced matching of size $\frac{1}{2}\binom{2r}{r}$, equality can be achieved, and there is a set $S$ such that $|S|=2r$ where $A_{i}\subseteq S$ and $B_{i}=S\setminus A_{i}$ for each $i$. Hence, one edge in a maximum induced matching $M$ uniquely determines the $2r$ elements of the ground set which can be used to form $M$.
Since a maximum induced matching $M$ can be uniquely determined by a single edge in $\textrm{KG}(n,r)$, this implies any two maximum induced matchings, i.e. matchings isomorphic to $\textrm{KG}(2r,r)$, are disjoint, and we can color each maximum induced matching bichromatically. Thus, we have a coloring of any $\textrm{KG}(n,r)$ with no monochromatic $\textrm{KG}(2r,r)$; i.e. $R^{\textrm{KG}}_{r,ind}(2r,2r)$ does not exist.
Finally, observe that for any $s,t\geq 2r$, $\textrm{KG}(s,r)$ and $\textrm{KG}(t,r)$ each contain a subgraph isomorphic to $\textrm{KG}(2r,r)$. Hence, since there exists a red/blue edge-coloring of $\textrm{KG}(n,r)$ with no induced monochromatic $\textrm{KG}(2r,r)$ for any $n\geq 2r$, in the same edge-coloring there are no monochromatic induced $\textrm{KG}(s,r)$ nor monochromatic induced $\textrm{KG}(t,r)$. ∎
Proof of Theorem 14.
Set $n=R_{r}(3r,3r)-1$. We aim to show that there is a red/blue coloring $c$ of $G=\textrm{KG}(n,r)$ containing no red induced $\textrm{KG}(3r,r)$ and no blue triangle. We begin by partitioning $V(G)$ into two sets. Consider a red/blue edge coloring $c^{\prime}$ of $G^{\prime}=K^{r}_{n}$ containing no monochromatic $K^{r}_{3r}$. Observe that $E(G^{\prime})=V(G)$ and there is a clear bijection between the two sets. Let us partition $V(G)$ into two sets $P_{1},P_{2}$ as follows: For $e=\{v_{1},v_{2},\dots,v_{r}\}\in E(G^{\prime})$, if $c(e)=R$, add $\{v_{1},v_{2},\dots,v_{r}\}$ to $P_{1}$. Otherwise, add $\{v_{1},v_{2},\dots,v_{r}\}$ to $P_{2}$.
Observe that by the choice of $c^{\prime}$, neither $G[P_{1}]$ nor $G[P_{2}]$ contain an induced copy of $\textrm{KG}(3r,r)$ (since otherwise, $c^{\prime}$ would contain a monochromatic $K^{r}_{3r}$). Therefore, for any edge $e$ within $G[P_{1}]$ or $G[P_{2}]$, set $c(e)=R$ and otherwise, set $c(e)=B$. Since the blue subgraph is a bipartite graph, it is triangle-free. And so, $c$ is a red/blue edge coloring of $G$ containing no red induced $\textrm{KG}(3r,r)$ and no blue triangle. ∎
5 Concluding remark
The multicolor Ramsey number $R(s_{1},s_{2},\dots,s_{t})$ is the minimum integer $n$ such that in any $t$-coloring of the edges of $K_{n}$, there exists a color-$i$ copy of $K_{s_{i}}$ for some $i\in[t]$. We can define the multicolor $r$-Kneser Ramsey number $R^{\textrm{KG}}_{r}(s_{1},\dots,s_{t})$ analogously.
While we do not consider multicolor Kneser Ramsey numbers in depth in this paper, by similar arguments as in the proof of Theorem 17, we obtain an upper bound on these numbers.
Theorem 20.
Fix $s_{1},s_{2},\dots,s_{t}\geq 3$ and for $1\leq i\leq t$, set $m_{i}=R(s_{1},s_{2},\dots,s_{i}-1,\dots,s_{t})$. Suppose $m_{1}\geq m_{j}$ for all $j\geq 2$. Then,
$R^{\textrm{KG}}_{r}(s_{1},\dots,s_{t})\leq(m_{1}+1)r+\sum_{i=2}^{t}(m_{i}-1).$
Appendix A: A good coloring of $\textrm{KG}(8,2)$.
Here, we provide a coloring of $\textrm{KG}(8,2)$ with no monochromatic triangle.
The red edges are:
((1,2),(3,4)),((1,2),(3,5)),((1,2),(3,6)),((1,2),(3,7)),((1,2),(3,8)),((1,3),(2,4)),((1,3),(2,5)),((1,3),(2,6)),((1,3),(2,7)),((1,3),(2,8)),((1,4),(2,3)),((1,4),(2,6)),((1,4),(2,8)),((1,4),(3,6)),((1,4),(3,8)),((1,4),(5,6)),((1,4),(6,8)),((1,5),(2,3)),((1,5),(2,6)),((1,5),(2,7)),((1,5),(3,6)),((1,5),(3,7)),((1,5),(6,7)),((1,5),(6,8)),((1,5),(7,8)),((1,6),(2,3)),((1,6),(2,4)),((1,6),(2,5)),((1,6),(3,4)),((1,6),(3,5)),((1,6),(4,5)),((1,6),(5,7)),((1,6),(5,8)),((1,7),(2,3)),((1,7),(2,5)),((1,7),(2,8)),((1,7),(3,5)),((1,7),(3,8)),((1,7),(4,8)),((1,7),(5,8)),((1,8),(2,3)),((1,8),(2,4)),((1,8),(2,7)),((1,8),(3,4)),((1,8),(3,7)),((1,8),(4,6)),((1,8),(4,7)),((1,8),(6,7)),((2,4),(3,6)),((2,4),(3,8)),((2,4),(5,6)),((2,4),(6,8)),((2,5),(3,6)),((2,5),(3,7)),((2,5),(6,7)),((2,5),(6,8)),((2,5),(7,8)),((2,6),(3,4)),((2,6),(3,5)),((2,6),(4,5)),((2,6),(5,7)),((2,6),(5,8)),((2,7),(3,5)),((2,7),(3,8)),((2,7),(4,5)),((2,7),(4,8)),((2,7),(5,8)),((2,8),(3,4)),((2,8),(3,7)),((2,8),(4,5)),((2,8),(4,6)),((2,8),(4,7)),((3,4),(5,6)),((3,4),(6,8)),((3,5),(6,7)),((3,5),(6,8)),((3,5),(7,8)),((3,6),(4,5)),((3,6),(5,7)),((3,6),(5,8)),((3,7),(4,8)),((3,7),(5,8)),((3,8),(4,6)),((3,8),(4,7)),((3,8),(6,7)),((4,5),(6,7)),((4,5),(6,8)),((4,5),(7,8)),((4,6),(5,7)),((4,6),(5,8)),((4,6),(7,8)),((4,7),(5,6)),((4,7),(5,8)),((4,7),(6,8)),((4,8),(5,6)),((4,8),(5,7)),((4,8),(6,7)),((5,6),(7,8)),((5,7),(6,8)),((5,8),(6,7))((1,2),(3,4)),((1,2),(3,5)),((1,2),(3,6)),((1,2),(3,7)),((1,2),(3,8)),((1,3),(2,4)),\\ ((1,3),(2,5)),((1,3),(2,6)),((1,3),(2,7)),((1,3),(2,8)),((1,4),(2,3)),((1,4),(2,6)),\\ ((1,4),(2,8)),((1,4),(3,6)),((1,4),(3,8)),((1,4),(5,6)),((1,4),(6,8)),((1,5),(2,3)),\\ ((1,5),(2,6)),((1,5),(2,7)),((1,5),(3,6)),((1,5),(3,7)),((1,5),(6,7)),((1,5),(6,8)),\\ ((1,5),(7,8)),((1,6),(2,3)),((1,6),(2,4)),((1,6),(2,5)),((1,6),(3,4)),((1,6),(3,5)),\\ ((1,6),(4,5)),((1,6),(5,7)),((1,6),(5,8)),((1,7),(2,3)),((1,7),(2,5)),((1,7),(2,8)),\\ ((1,7),(3,5)),((1,7),(3,8)),((1,7),(4,8)),((1,7),(5,8)),((1,8),(2,3)),((1,8),(2,4)),\\ ((1,8),(2,7)),((1,8),(3,4)),((1,8),(3,7)),((1,8),(4,6)),((1,8),(4,7)),((1,8),(6,7)),\\ ((2,4),(3,6)),((2,4),(3,8)),((2,4),(5,6)),((2,4),(6,8)),((2,5),(3,6)),((2,5),(3,7)),\\ ((2,5),(6,7)),((2,5),(6,8)),((2,5),(7,8)),((2,6),(3,4)),((2,6),(3,5)),((2,6),(4,5)),\\ ((2,6),(5,7)),((2,6),(5,8)),((2,7),(3,5)),((2,7),(3,8)),((2,7),(4,5)),((2,7),(4,8)),\\ ((2,7),(5,8)),((2,8),(3,4)),((2,8),(3,7)),((2,8),(4,5)),((2,8),(4,6)),((2,8),(4,7)),\\ ((3,4),(5,6)),((3,4),(6,8)),((3,5),(6,7)),((3,5),(6,8)),((3,5),(7,8)),((3,6),(4,5)),\\ ((3,6),(5,7)),((3,6),(5,8)),((3,7),(4,8)),((3,7),(5,8)),((3,8),(4,6)),((3,8),(4,7)),\\ ((3,8),(6,7)),((4,5),(6,7)),((4,5),(6,8)),((4,5),(7,8)),((4,6),(5,7)),((4,6),(5,8)),\\ ((4,6),(7,8)),((4,7),(5,6)),((4,7),(5,8)),((4,7),(6,8)),((4,8),(5,6)),((4,8),(5,7)),\\ ((4,8),(6,7)),((5,6),(7,8)),((5,7),(6,8)),((5,8),(6,7))
The remaining edges are blue.
We can describe part of this coloring as follows. First, we let $A=\binom{[3]}{2}$, $B=\binom{[8]-[3]}{2}$, and $C=V(G)-(A\cup B)$. Then all the edges from $A$ to $C$ as well as all the edges within $B$ are red, and all the edges from $A$ to $B$ are blue.
Appendix B: Our code
Below is the code we used to obtain the results in Theorem 12.
    import itertools as it
    from pysat.solvers import Glucose3
    Builds Kneser Graph Kn(n, r)
    Input: n, r
    Output: (V, E)
        V (list): Vertices (all r-subsets of [n]) as r-tuples
        E (list of 2-tuples): Edges as a pair of r-tuples
    def build_kneser_graph(n,r):
        V = list(it.combinations(range(1, n + 1), r))
        E = []
        for (a, b) in it.combinations(V, 2):
            if set(a).isdisjoint(set(b)):
                E.append((a, b))
        return (V, E)
    Try to find if Kn(n,r) can be 2-colored so that there is no red K_s or blue K_t
    For no red K_s, for every s-clique, need at least 1 blue edge.
    That is, if e_1, ..., e_k are the edges of an s-clique and e_i = 1 if red, e_i = -1 if blue,
    want to satisfy (-e_1 v -e_2 v ... v -e_k)
    For no blue K_t, for every t-clique, need at least 1 red edge.
    That is, if e_1, ..., e_k are the edges of a t-clique and e_i = 1 if red, e_i = -1 if blue,
    want to satisfy (-e_1 v -e_2 v ... v -e_k)
    (NOTE: 1 corresponds to True, -1 corresponds to False)
    If there is a good coloring, outputs (True, R, B), where R = red edges, B = blue edges
    If there is no good coloring, outputs (False, [], [])
    def build_sat_instances(n, r, s, t):
        solver = Glucose3()
        clauses = []
        V = []
        E = []
        #Create vertex and edge list
        (V, E) = build_kneser_graph(n, r)
        for s_tuple in it.combinations(V, s):
            #check that the s-tuple forms a clique in Kn(n,r)
            if all([set(s_tuple[i]).isdisjoint(set(s_tuple[j])) for (i,j) in it.combinations(range(len(s_tuple)), 2)]):
                #Add clause forbidding monochromatic K_s and K_t (if s = t)
                if s == t:
                    cliq = []
                    for edg in it.combinations(s_tuple, 2):
                        idx = -1
                        try:
                            idx = E.index(edg)
                        except:
                            idx = E.index((edg[1], edg[0]))
                        cliq.append(idx + 1)
                    solver.add_clause([-x for x in cliq])
                    solver.add_clause(cliq)
                #Add clause forbidding red K_s
                if s != t:
                    cliq = []
                    for edg in it.combinations(s_tuple, 2):
                        idx = -1
                        try:
                            idx = E.index(edg)
                        except:
                            idx = E.index((edg[1], edg[0]))
                        cliq.append(idx + 1)
                    solver.add_clause([-x for x in cliq])
        if s != t:
            for t_tuple in it.combinations(V, t):
                #check that the t-tuple forms a clique in Kn(n,r)
                if all([set(t_tuple[i]).isdisjoint(set(t_tuple[j])) for (i,j) in it.combinations(range(len(t_tuple)), 2)]):
                    #Add clause forbidding red K_t
                    cliq = []
                    for edg in it.combinations(t_tuple, 2):
                        idx = -1
                        try:
                            idx = E.index(edg)
                        except:
                            idx = E.index((edg[1], edg[0]))
                        cliq.append(idx + 1)
                    solver.add_clause(cliq)
        #Solve SAT problem and recover coloring
        result = solver.solve()
        R = []
        B = []
        if result:
            L = solver.get_model()
            R = [E[i - 1] for i in L if i > 0]
            B = [E[-i - 1] for i in L if i < 0]
        return (result, R, B)
    #output: True, R, B
    build_sat_instances(8, 2, 3, 3)
    #output: False, [], []
    build_sat_instances(9, 2, 3, 3)
    #output: False, [], []
    build_sat_instances(13, 3, 3, 3)

### v2

@@
 It is easy to see that $\beta\leq 3$.
-Indeed, let $S_{1},\dots,S_{6}\in{[3k]\choose k}$ be six sets with the property that for every $\in[3k], tt$ belongs to exactly two of the sets $S_{1},\dots,S_{6}$.
+Indeed, let $S_{1},\dots,S_{6}\in{[3k]\choose k}$ be six sets with the property that for every $\in[3k], ii$ belongs to exactly two of the sets $S_{1},\dots,S_{6}$.
 (It is easy to construct such sets $S_{1},\dots,S_{6}$: take the six distinct sets in two disjoint perfect matchings in the $k$-uniform hypergraph on vertex set $[3k]$.) Now consider a red/blue edge-coloring of the complete graph on vertices $S_{1},\dots,S_{6}$.
 Since $R(3,3)=6$, this graph contains a monochromatic triangle $S_{i}S_{j}S_{k}$, and $S_{i}\cap S_{j}\cap S_{k}=\emptyset$ for every distinct $i,j,k\in[6]$.
-In this paper we make progress on Problems 2 and 4 in two different ways. First, in Section 2 we improve the upper bound on $\beta$ to $7/3$.
+In this paper we make progress on Problems 2-4 in two different ways. First, in Section 2 we improve the upper bound on $\beta$ in Problem 4 to $7/3$.
 Theorem 5.
@@
 Theorem 6.
-For any $0<\alpha<\frac{1}{4}$, there exists some large enough $n$ such that any red/blue edge-coloring of $\textrm{KG}(n,\alpha n)$ contains a monochromatic triangle.
-Note that for $\alpha=\frac{1}{4}$, for $n=8$ there exists a red/blue edge-coloring of $\textrm{KG}(n,\alpha n)$ with no monochromatic triangle, as implied by Theorem 12.
+For any rational number $\alpha$ such $0<\alpha<\frac{1}{4}$, there exists some large enough integer $n$ such that $\alpha n$ is an integer and any red/blue edge-coloring of $\textrm{KG}(n,\alpha n)$ contains a monochromatic triangle.
+Note that when $\alpha=\frac{1}{4}$, for $n=8$ there exists a red/blue edge-coloring of $\textrm{KG}(n,\alpha n)$ with no monochromatic triangle, as implied by Theorem 12 (an explicit coloring is given in Appendix A).
+We do not know if this holds for larger $n$.
 We then continue with a systematic study of Kneser Ramsey numbers for various parameters $r,s,t$. The following trivial upper bound on the $r$-Kneser Ramsey number follows from Ramsey’s theorem:
@@
 Theorem 7.
-Fix $s,t\geq 3$ and suppose $R(s-1,t)\geq R(s,t-1)$. Set M=R(s1,t)M=R(s-1,t) and m=R(s,t1)m=R(s,t-1). We have
-$R^{\textrm{KG}}_{r}(s,t)\leq(M+1)r+m-1.$
-In particular,
+Fix $s,t\geq 3$ and suppose $R(s-1,t)\geq R(s,t-1)$. Set M=R(s1,t)M=R(s-1,t) and m=R(s,t1)m=R(s,t-1). Then
+$R^{\textrm{KG}}_{r}(s,t)\leq(M+1)r+m-1,$
+and in particular,
 $R^{\textrm{KG}}_{r}(s,s)\leq(R(s,s-1)+1)r+R(s,s-1)-1.$
-In Section 3 we establish improved upper bounds on $R^{\textrm{KG}}_{r}(s,t)$ for certain families of $s,t$.
+In Section 3 we establish further upper bounds on $R^{\textrm{KG}}_{r}(s,t)$ for certain families of $s,t$.
 Theorem 8.
@@
 Theorem 9.
-For $t\geq 5$, we have $R^{\textrm{KG}}_{2}(3,t)\leq 2(R(3,t-2))+2t+1.$
+For $t\geq 5$, we have $R^{\textrm{KG}}_{r}(3,t)\leq rR(3,t-2)+2r+2t-3.$
 We then derive a general lower bound, and an improved lower bound for $s=t=3$ and $r\geq 3$.
@@
 For $r\geq 2$, we have $R^{\textrm{KG}}_{r}(3,3)\geq 3r+2.$
-Using a computer code, we also obtain an exact value on $R^{\textrm{KG}}_{2}(3,3)$. The computational proof of the next theorem is discussed in Subsection 3.3.
+Using a computer code, we also obtain an exact value on $R^{\textrm{KG}}_{2}(3,3)$ and improve bounds on other Kneser Ramsey numbers.
+The computational proof of the next theorem is discussed in Subsection 3.3.
 Theorem 12.
-We have $R^{\textrm{KG}}_{2}(3,3)=9$ and $R^{\textrm{KG}}_{3}(3,3)\leq 13$.
+We have
+$R^{\textrm{KG}}_{2}(3,3)=9,~~R^{\textrm{KG}}_{3}(3,3)\leq 13,\text{~~and~~}12\leq R^{\textrm{KG}}_{2}(4,3)\leq 14.$
 Some of our results are summarized in Table 1 (for $r=2$) and Table 2 (for $r=3$).
@@
 The subscripts on the numbers are the theorems in which we obtain the bound.
-For example, we have $11\leq R^{\textrm{KG}}_{2}(3,4)\leq 16$, where the lower bound follows from Theorem 10 and the upper bound is proven in Theorem 8.
+For example, we have $16\leq R^{\textrm{KG}}_{2}(3,5)\leq 23$, where the lower bound follows from Theorem 10 and the upper bound is proven in Theorem 9.
 If there is only one number, then it is an exact value.
@@
 $s$ $t$ $t=3$ $t=4$ $t=5$ $t=6$
-$s=3$ 912 1110, 168 1610, 239 2010, 319
+$s=3$ 912 1212, 1412 1610, 239 2010, 319
 $s=4$ 2010, 287 2710, 488 3810, 697
-$s=5$ 4510, 767 6010, 1348
+$s=5$ 4510, 767 6010, 1328
 $s=6$ 10410, 2627
@@
 $s$ $t$ $t=3$ $t=4$ $t=5$ $t=6$
-$s=3$ 1111, 1312 1310, 247 1810, 347 2210, 507
-$s=4$ 2210, 387 2910, 707 4010, 957
-$s=5$ 4710, 1027 6210, 1877
+$s=3$ 1111, 1312 1310, 228 1810, 319 2210, 429
+$s=4$ 2210, 387 2910, 648 4010, 957
+$s=5$ 4710, 1027 6210, 1748
 $s=6$ 10610, 3507
 We conclude this paper with some progress on a Ramsey-type question that was posted by Dömötör Pálvölgyi on mathoverflow in 2016 [6].
-Here an induced $\textrm{KG}(3r,r)$ is a subgraph of $\textrm{KG}(n,r)$ induced on vertex set $S\choose r$}, where $S\subset[n]$ is a set of size $3r$.
+Given integer $n\geq s\geq r$, an induced $\textrm{KG}(s,r)$ is a subgraph $G$ of the graph $\textrm{KG}(n,r)$ induced on a vertex set $S\choose r$}, where $S\subset[n]$ is a set of size $s$.
+We say that $S$ is the ground set of $G$.
 Problem 13 (Pálvölgyi [6]).
@@
 Theorem 15 (Alon - Frankl - Lovász [1]).
-Let $k_{1}\geq k_{2}\geq\dots\geq k_{t}\geq 2$ and suppose $n\geq k_{1}r+\sum_{i=2}^{t}(k_{i}-1)$.
+Let $n\geq 2r\geq 2$ and $k_{1}\geq k_{2}\geq\dots\geq k_{t}\geq 2$ be integers, and suppose $n\geq k_{1}r+\sum_{i=2}^{t}(k_{i}-1)$.
 Let $F_{1},F_{2},\dots,F_{t}$ be a partition of $\binom{[n]}{r}$.
 Then, for some $i\in[t]$, $F_{i}$ contains $k_{i}$ pairwise disjoint sets.
-Observe that when $k_{1}=k_{2}=\cdots=k_{t}=2$, Theorem 15 recovers the chromatic number of the Kneser graph. When $t=2$, we apply a simpler form.
-Theorem 16 (Alon - Frankl - Lovász [1]).
-Let $n\geq(r+1)k-1$ and let $F_{1},F_{2}$ be a partition of $\binom{[n]}{r}$. Then, either $F_{1}$ or $F_{2}$ contains $k$ pairwise disjoint sets.
+Observe that when $k_{1}=k_{2}=\cdots=k_{t}=2$, Theorem 15 recovers the chromatic number of the Kneser graph.
 2 The Holmsen - Hrusak - Roldán-Pensado problem
@@
 Finally, if the triangle $H_{1}H_{2}H_{3}$ is monochromatic red, we have a symmetric argument by considering instead the induced graph on vertex set $\{F_{1},F_{3},H_{1},H_{2},H_{3}\}$. ∎
-Our next goal is to prove Theorem 6, which gives a positive answer to Problem 2 in the case $\alpha<1/4$ and $t=3$. In order to prove the theorem, we need the following auxiliary result.
+Our next goal is to prove Theorem 6, which gives a positive answer in Problem 2 in the case $\alpha<1/4$ and $t=3$. In order to prove the theorem, we need the following auxiliary result.
 Theorem 16.
@@
 Set M=R(s1,t)M=R(s-1,t) and m=R(s,t1)m=R(s,t-1).
-Let $0<\varepsilon<\frac{1}{M+1}$ and set $\alpha=\frac{1}{M+1}-\varepsilon$.
-Then, there exists some $n\in\mathbb{N}$ such that $\alpha n\in\mathbb{N}$ and $\textrm{KG}(n,\alpha n)$ contains a red $K_{s}$ or a blue $K_{t}$.
-Proof of Theorem 17.
-Fix $r\geq 2$ and set Nr=Mr+m1N-r=Mr+m-1.
-First, we will show that in any red/blue coloring of $G=\textrm{KG}(N,r)$, there is a red $K_{s}$ or a blue $K_{t}$.
-Suppose for contradiction that there is some coloring $c:E(G)\to\{R,B\}$ of $G$ without a red $K_{s}$ or blue $K_{t}$.
+Let $\alpha$ be a rational number so that $0<\alpha<\frac{1}{M+1}$.
+Then, there exists some integer $n$ such that $\alpha n$ is an integer and $\textrm{KG}(n,\alpha n)$ contains a red $K_{s}$ or a blue $K_{t}$.
+Proof.
+Fix an integer $r\geq 2$ and let N=r(M+1)+m1N=r(M+1)+m-1.
+We first show that in any red/blue coloring of $G=\textrm{KG}(N,r)$, there is a red $K_{s}$ or a blue $K_{t}$.
+Indeed, assume for contradiction that there is some coloring $c:E(G)\to\{R,B\}$ of $G$ with no red $K_{s}$ or blue $K_{t}$.
 Let $v=[N]\setminus[N-r]\in V(G)$.
-Then, $\binom{[N-r]}{r}=\binom{[Mr+m-1]}{r}$ can be partitioned into $N_{R}(v)$ and $N_{B}(v)$.
-Since $c$ contains no red $K_{s}$ and no blue $K_{t}$, then $G[N_{R}(v)]$ contains no red $K_{s-1}$ and no blue $K_{t}$.
-Similarly, $G[N_{B}(v)]$ contains no red $K_{s}$ and no blue $K_{t-1}$.
+Then, $\binom{[N-r]}{r}=\binom{[Mr+m-1]}{r}$ can be partitioned as $N_{R}(v)\cup N_{B}(v)$, where $N_{R}(v)$ and $N_{B}(v)$ are the red and blue neighborhoods of $v$, respectively.
+Since $c$ contains no red $K_{s}$ and no blue $K_{t}$, then $G[N_{R}(v)]$ contains no red $K_{s-1}$ and no blue $K_{t}$, and similarly, $G[N_{B}(v)]$ contains no red $K_{s}$ and no blue $K_{t-1}$.
 In particular, $G[N_{R}(v)]$ contains no $K_{M}$ and $G[N_{B}(v)]$ contains no $K_{m}$.
-However, by Theorem 16, setting $F_{1}=N_{R}(v)$ and $F_{2}=N_{B}(v)$, either $N_{R}(v)$ contains a clique of size $M$ or $N_{B}(v)$ contains a clique of size $m$, a contradiction.
-For the general case, fix $0<\varepsilon<\frac{1}{M+1}$ and set $\alpha=\frac{1}{M+1}-\varepsilon$.
-Choose $r\in\mathbb{Z}$ such that $\frac{r}{(M+1)r+m-1}\geq\alpha$.
-Set Nr=Mr+m1N-r=Mr+m-1 and $n=\frac{r}{\alpha}$.
-Then, by choice of $r$, observe that $N\leq n$.
-Therefore, since $\textrm{KG}(N,\alpha n)\subseteq\textrm{KG}(n,\alpha n)$, any red-blue coloring of $\textrm{KG}(n,\alpha n)$ contains a red $K_{s}$ or a blue $K_{t}$.
+However, by Theorem 15, setting $F_{1}=N_{R}(v)$ and $F_{2}=N_{B}(v)$, we have that either $N_{R}(v)$ contains a clique of size $M$ or $N_{B}(v)$ contains a clique of size $m$, a contradiction.
+Now, let $\alpha$ be a rational number so that $0<\alpha<\frac{1}{M+1}$.
+Let $r$ be a positive integer so that $\frac{r}{(M+1)r+m-1}\geq\alpha$ and $\frac{r}{\alpha}$ is an integer.
+Let N=r+Mr+m1N=r+Mr+m-1 and $n=\frac{r}{\alpha}$.
+By the choice of $r$ we have $n\geq N$.
+Therefore, $\textrm{KG}(N,r)\subseteq\textrm{KG}(n,\alpha n)$, and thus any red-blue coloring of $\textrm{KG}(n,\alpha n)$ contains a red $K_{s}$ or a blue $K_{t}$.
 ∎
-Note that the general bound on $R^{\textrm{KG}}_{r}(s,t)$ in Theorem 7 is proven in the first paragraph of the proof of Theorem 17 above.
+Note that the general bound on $R^{\textrm{KG}}_{r}(s,t)$ in Theorem 7 is proven in the first paragraph of the proof of Theorem 16.
 We are now ready to prove Theorem 6.
@@
 This gives $M=m=R(2,3)=3$.
-For any $0<\varepsilon<\frac{1}{4}$, we have $\alpha=\frac{1}{4}-\varepsilon$, and by Theorem 17 there exists some $n\in\mathbb{N}$ with $\alpha n\in\mathbb{N}$ such that any red/blue coloring of the edges of $\textrm{KG}(n,\alpha n)$ contains a monochromatic triangle.
+Let $\alpha$ be a rational number satisfying $0<\alpha<\frac{1}{4}$.
+Then $\alpha<\frac{1}{M+1}$, and thus by Theorem 16 there exists some integer $n$ such that $\alpha n$ is an integer and any red/blue coloring of the edges of $\textrm{KG}(n,\alpha n)$ contains a monochromatic triangle.
 ∎
@@
 3.1 Upper bounds
+The proofs of both Theorem 8 and Theorem 9 follow the same scheme, described as follows.
+For some integers $N,r,s,t$, let $G=\textrm{KG}(N,r)$ and let $c:E(G)\to\{R,B\}$ be a red/blue edge-coloring of $G$.
+Assume that $c$ has no red $K_{s}$ or blue $K_{t}$, and suppose there exists a blue edge in $c$.
+Then (by renaming the integers in $[N]$) we may assume without loss of generality that $c(uv)=B$, where $u=\{N,N-1,\dots,N-r+1\}\in V(G)$ and $v=\{N-r,N-r-1,\dots,N-2r+1\}\in V(G)$.
+Note that $G[N_{G}(u)]\cap G[N_{G}(v)]=\textrm{KG}(N-2r,r)$.
+We now partition the set of vertices $N-2r]\choose r$} of $G^{\prime}=\textrm{KG}(N-2r,r)$ into four sets $N_{XY}$, where $X,Y\in\{R,B\}$, where $N_{XY}$ is the set of vertices $w\in{[N-2r]\choose r}$ such that $c(wu)=X$ and $c(wv)=Y$.
+Observe that $N_{BR}$ contains no red $K_{s-1}$, for otherwise we have a red $K_{s}$ together with $v$, and additionally, $N_{BR}$ contains no blue $K_{t-1}$, for otherwise have a blue $K_{t}$ together with $u$.
+Thus
+$\omega(G^{\prime}[N_{BR}])\leq R(s-1,t-1)-1$
+where $\omega(H)$ is the clique number of the graph $H$. Similarly, $N_{RR}\cup N_{RB}$ contains no red $K_{s-1}$ or blue $K_{t}$, so
+$\omega(G^{\prime}[N_{RR}\cup N_{RB}])\leq R(s-1,t)-1.$
+Finally $N_{BB}$ contains no red $K_{s}$ and no blue $K_{t-2}$, as otherwise we can find a blue $K_{t}$ together with the edge $uv$. Hence
+$\omega(G^{\prime}[N_{BB}])\leq R(s,t-2)-1.$
+Now we would like to apply Theorem 15 to conclude that at least one of the above three induced subgraph of $G^{\prime}$ contains a large clique, which will constitute a contradiction.
+But in order to do that, we need to compare the sizes of the Ramsey numbers R(s1,t1)R(s-1,t-1), R(s1,t)R(s-1,t), and R(s,t2)R(s,t-2), which we do not know how to do in general.
+However, in the two special cases where $s$ and $t$ differ by 1 (Theorem 8) or when $s=3$ (Theorem 9) we can find the largest value among the three Ramsey numbers, which allows us to apply Theorem 15.
 Proof of Theorem 8.
@@
 $R^{\textrm{KG}}_{r}(s-1,s)\leq rR(s-2,s)+2(R(s-2,s-1)-1)+2r.$
-Let N=rR(s2,s)+2R(s2,s1)+2rN=rR(s-2,s)+2R(s-2,s-1)+2r, $G=\textrm{KG}(N,r)$ and let $c:E(G)\to\{R,B\}$ be a red/blue edge-coloring of $G$.
-For sake of contradiction, suppose there is no red $K_{s-1}$ nor blue $K_{s}$.
-If there is no blue edge in $G$, then we are done.
-So, set $u=\{N,N-1,\dots,N-r+1\}\in V(G)$ and $v=\{N-r,N-r-1,\dots,N-2r+1\}\in V(G)$.
-Without loss of generality, suppose $c(u,v)=B$.
-Then, $N(u)\cap N(v)\cong G^{\prime}=\textrm{KG}(N-2r,r)$.
-We will partition the vertices of $G^{\prime}$ into 4 sets.
-For $(X,Y)\in\{(R,R),(R,B),(B,R),(B,B)\}$, we denote by $N_{XY}$ the set of vertices $w$ in $G^{\prime}$ such that $c(w,u)=X$ and $c(w,v)=Y$.
-Observe that $N_{BR}$ must contain no red $K_{s-2}$ or we may find a red $K_{s-1}$ along with $v$, and $N_{BR}$ must contain no blue $K_{s-1}$ or we may find a blue $K_{s}$ along with $u$.
-Thus $\omega(G^{\prime}[N_{BR}])\leq R(s-2,s-1)-1$.
-Similarly, $N_{RR}\cup N_{RB}$ must contain no red $K_{s-2}$ or blue $K_{s}$, so $\omega(G^{\prime}[N_{RR}\cup N_{RB}])\leq R(s-2,s)-1$.
-Furthermore, $N_{BB}$ must contain no red $K_{s-1}$ and no blue $K_{s-2}$, as otherwise we can find a blue $K_{s}$ along with the edge $uv$.
-Hence $\omega(G^{\prime}[N_{BB}])\leq R(s-2,s-1)-1$.
-However, invoking Theorem 15 with $F_{1}=N_{RR}\cup N_{RB}$ ($k_{1}=R(s-2,s)$), $F_{2}=N_{BR}$ ($k_{2}=R(s-2,s-1)$), and $F_{3}=N_{BB}$ ($k_{3}=R(s-2,s-1)$), since $k_{1}\cdot r+(k_{2}-1)+(k_{3}-1)=N-2r$, either $F_{1}$ contains a clique on R(s2,s)R(s-2,s) vertices, $F_{2}$ contains a clique on R(s2,s1)R(s-2,s-1) vertices, or $F_{3}$ contains a clique on R(s2,s1)R(s-2,s-1) vertices.
-In any case, we are guaranteed either a red $K_{s-1}$ or a blue $K_{s}$, a contradiction.
-Therefore, $R^{\textrm{KG}}_{r}(s-1,s)\leq N$.
+Let N=rR(s2,s)+2(R(s2,s1)1)+2rN=rR(s-2,s)+2(R(s-2,s-1)-1)+2r, $G=\textrm{KG}(N,r)$ and let $c:E(G)\to\{R,B\}$ be a red/blue edge-coloring of $G$.
+Assume for contradiction that $c$ has no red $K_{s-1}$ or blue $K_{s}$.
+Since
+$|V(G)|={N\choose r}\geq N>rR(s-2,s)=rR(s,s-2),$
+there is a clique of size R(s,s2)R(s,s-2) in $G$, so if $c$ has no blue edge it must have a red $K_{s}$.
+So we may assume some edge is colored blue.
+Without loss of generality, suppose $c(uv)=B$, where $u=\{N,N-1,\dots,N-r+1\}\in V(G)$ and $v=\{N-r,N-r-1,\dots,N-2r+1\}\in V(G)$.
+Then $G[N_{G}(u)]\cap G[N_{G}(v)]=\textrm{KG}(N-2r,r)$.
+Let $G^{\prime}=\textrm{KG}(N-2r,r)$.
+As explained above, we have (for $(s,t)\leftarrow(s-1,s)$)
+$\begin{split}\omega(G^{\prime}[N_{BR}])&\leq R(s-2,s-1)-1,\\ \omega(G^{\prime}[N_{RR}\cup N_{RB}])&\leq R(s-2,s)-1,\\ \omega(G^{\prime}[N_{BB}])&\leq R(s-1,s-2)-1=R(s-2,s-1)-1.\end{split}$ (1)
+Clearly, we have $R(s-2,s)\geq R(s-2,s-1)$, and thus we can invoke Theorem 15 with $F_{1}=N_{RR}\cup N_{RB}$, $F_{2}=N_{BR}$ and $F_{3}=N_{BB}$, and with $k_{1}=R(s-2,s),$ $k_{2}=R(s-2,s-1),$ $k_{3}=R(s-2,s-1)$.
+Since
+$k_{1}\cdot r+(k_{2}-1)+(k_{3}-1)=N-2r,$
+Theorem 15 entails that either $F_{1}$ contains a clique on R(s2,s)R(s-2,s) vertices, or $F_{2}$ contains a clique on R(s2,s1)R(s-2,s-1) vertices, or $F_{3}$ contains a clique on R(s2,s1)R(s-2,s-1) vertices, contradicting (1).
 ∎
-Theorem 9 gives an improvement when $r=2$ and $s=3$.
 Proof of Theorem 9.
-For sake of contradiction, fix N>2(R(3,t2))+2tN>2(R(3,t-2))+2t, let $G=\textrm{KG}(N,2)$, $c:E(G)\to\{R,B\}$ be a red/blue edge-coloring of $G$, and suppose there is no red $K_{3}$ or blue $K_{t}$.
-Note that if there are no blue edges, then there must be a red $K_{3}$, and we have arrived at our contradiction.
-Hence, without loss of generality, we may assume $c(\{1,2\},\{3,4\})=B$.
-Next, we look at $N(\{1,2\})\cap N(\{3,4\})\cong G^{\prime}=\textrm{KG}(N-4,2)$, and we will partition the vertices of $G^{\prime}$ into $4$ sets.
-For $(X,Y)\in\{(R,R),(R,B),(B,R),(B,B)\}$, we denote by $N_{XY}$ the set of vertices $v$ in $G^{\prime}$ such that $c(v,\{1,2\})=X$ and $c(v,\{3,4\})=Y$.
-First, observe that for any two vertices $u,v\in N_{RB}$ with $uv\in E(G^{\prime})$, $c(u,v)=B$ (otherwise there is a red $K_{3}$ with $u,v,\{1,2\}$, a contradiction).
-Similarly, for each case of $u,v\in N_{RR}$, $u,v\in N_{BR}$, and $u\in N_{RR},v\in N_{RB}$, we have $c(u,v)=B$.
-These observations allow us to bound the clique numbers of $N_{RR}\cup N_{RB}$ and $N_{BR}$.
-In particular, $\omega(G^{\prime}[N_{RR}\cup N_{RB}])\leq t-1$ since all edges are blue and we cannot have a blue $K_{t}$.
-Additionally, $\omega(G^{\prime}[N_{BR}])\leq t-2$ since all edges are blue, and if there were a blue $K_{t-1}$, together with vertex $\{1,2\}$, we would have a blue $K_{t}$.
-Finally, $\omega(G^{\prime}[N_{BB}])<R(t-2,3)$ since otherwise, there would be either a copy of a red $K_{3}$ (a contradiction), or a blue copy of $K_{t-2}$, which together with the blue edge $(\{1,2\},\{3,4\})$ would be a blue $K_{t}$, another contradiction.
-By Theorem 15, with $F_{1}=N_{RR}\cup N_{RB}$ ($k_{1}=R(t-2,3))$, $F_{2}=N_{BR}$ ($k_{2}=t$), and $F_{3}=N_{BB}$ ($k_{3}=t-1$), since $N-4\geq 2(R(t-2,3))+(t-1)+(t-1-1)$, either $F_{1}$ contains a $K_{R(t-2,3)}$, $F_{2}$ contains a $K_{t}$, or $F_{3}$ contains a $K_{t-1}$, a contradiction in all cases.
-This final contradiction proves the result.
-∎
+We now apply our proof scheme with $(s,t)\leftarrow(3,t)$. We would like to show that for $t\geq 5$,
+$R^{\textrm{KG}}_{r}(3,t)\leq rR(3,t-2)+2r+2t-3.$
+Let N=rR(3,t2)+2r+2t3N=rR(3,t-2)+2r+2t-3 and $G=\textrm{KG}(N,r)$.
+Let $c:E(G)\to\{R,B\}$ be a red/blue edge-coloring of $G$.
+Assume for contradiction that $c$ has no red $K_{3}$ nor blue $K_{t}$.
+Like before, since N>rR(3,t2)N>rR(3,t-2), if $c$ has no blue edge, then it must have a red $K_{3}$.
+So we may assume some edge is colored blue, and without loss of generality, suppose $c(uv)=B$, where $u$ and $v$ as before.
+Then $G[N_{G}(u)]\cap G[N_{G}(v)]=\textrm{KG}(N-2r,r)$.
+Let $G^{\prime}=\textrm{KG}(N-2r,r)$.
+Now we have
+$\begin{split}\omega(G^{\prime}[N_{BR}])&\leq R(2,t-1)-1=t-2,\\ \omega(G^{\prime}[N_{RR}\cup N_{RB}])&\leq R(2,t)-1=t-1,\\ \omega(G^{\prime}[N_{BB}])&\leq R(3,t-2)-1.\end{split}$ (2)
+It is straightforward to see that R(3,t2)>2(t3)R(3,t-2)>2(t-3) (indeed, take two disjoint cliques on t3t-3 vertices each, and color their edges blue; color the edges between them red; then the red graph is bipartite so contains no triangles).
+This implies that for $t\geq 5$ we have $R(3,t-2)\geq t$.
+Therefore, applying Theorem 15 with $F_{1}=N_{RR}\cup N_{RB}$, $k_{1}=R(3,t-2)$, $F_{2}=N_{BR}$, $k_{2}=t$, $F_{3}=N_{BB}$, and $k_{3}=t-1$, we have
+$N-2r\geq r(R(3,t-2))+(t-1)+(t-1-1),$
+implying that for some $i\in[3]$, $F_{i}$ contains a clique of size $k_{i}$, contradicting (2). ∎
 3.2 Lower bounds
-The lower bound in Proposition 10 takes advantage of the chromatic number of Kneser graphs.
+For the lower bound in Proposition 10 we employ the chromatic number of Kneser graphs.
+It is easy to verify that $\chi(\textrm{KG}(n,r))\geq n-2r+2$ by exhibiting a explicit proper coloring, and this is the bound we need for the proof.
 Lovász [5] proved that $\chi(\textrm{KG}(n,r))=n-2r+2$, which was one of the earliest applications of topological methods to combinatorics.
 Proof of Proposition 10.
-Let $n=R^{\textrm{KG}}_{r}(s,t)$ and set $G=\textrm{KG}(n,r)$.
-Suppose $\chi(G)=n-2r+2<R(s,t)$ and let $c:V(G)\to[n-2r+2]$ be a proper vertex-coloring of $G$.
-Consider the complete graph $G^{\prime}$ on n2r+2n-2r+2 vertices with $V(G^{\prime})=[n-2r+2]$ and let $C^{\prime}:E(G^{\prime})\to\{R,B\}$ be a red/blue edge-coloring of $G^{\prime}$ avoiding a red copy of $K_{s}$ and a blue copy of $K_{t}$.
-Now, consider the following red/blue edge-coloring of $G$: For $uv\in E(G)$, color $uv$ with $C^{\prime}(c(u),c(v))$.
-We claim that this coloring has no red copy of $K_{s}$ or blue copy of $K_{t}$.
-Suppose there is a red copy of $K_{s}$ with vertices $\{v_{1},v_{2},\dots,v_{s}\}$.
-Observe that for all $i\neq j$, $c(v_{i})\neq c(v_{j})$ since $c$ is a proper vertex-coloring and $v_{i}v_{j}\in E(G)$.
-But now, the color of $v_{i}v_{j}$ is $C^{\prime}(c(v_{i}),c(v_{j}))=R$.
-This implies that the graph induced by the vertices $\{c(v_{1}),c(v_{2}),\dots,c(v_{s})\}$ is a red copy of $K_{s}$, a contradiction to our choice of $C^{\prime}$.
-Therefore, $n-2r+2\geq R(s,t)$.
+For $r\geq 1$ and $s,t\geq 2$, we wish to show $R^{\textrm{KG}}_{r}(s,t)\geq R(s,t)+2r-2.$ Let $n=R^{\textrm{KG}}_{r}(s,t)$ and set $G=\textrm{KG}(n,r)$.
+Assume for contradiction that $R(s,t)>n-2r+2=\chi(G)$ and let $c:V(G)\to[n-2r+2]$ be a proper vertex-coloring of $G$.
+Consider the complete graph $G^{\prime}$ on vertex set [n2r+2][n-2r+2] and let $c^{\prime}:E(G^{\prime})\to\{R,B\}$ be a red/blue edge-coloring of $G^{\prime}$ avoiding a red copy of $K_{s}$ and a blue copy of $K_{t}$.
+Note that $c^{\prime}$ exists because of our negation assumption.
+Now, define a red/blue edge-coloring of $f:E(G)\to\{R,B\}$ by
+$f(uv)=c^{\prime}(c(u),c(v))$
+for $uv\in E(G)$. We claim that $f$ has no red copy of $K_{s}$ or blue copy of $K_{t}$. Suppose there is a red copy of $K_{s}$ with vertices $\{v_{1},v_{2},\dots,v_{s}\}$. Then for $i\neq j$ we have
+$f(v_{i}v_{j})=c^{\prime}(c(v_{i}),c(v_{j}))=R.$
+Observe that when $i\neq j$, $c(v_{i})\neq c(v_{j})$ because $c$ is a proper vertex-coloring of $G$ and $v_{i}v_{j}\in E(G)$.
+This implies that the subgraph of $G^{\prime}$ induced on the vertices $\{c(v_{1}),c(v_{2}),\dots,c(v_{s})\}$ is a red copy of $K_{s}$, a contradiction to our choice of $c^{\prime}$.
+A similar argument shows that $f$ has no blue copy of $K_{t}$.
 ∎
@@
 $Z$ $=V(G)-Y-X\simeq V(\textrm{KG}(3r-1,r)).$
-Observe that the induced subgraph $G[Z]$ of $G$ on $Z$ is triangle-free.
+Observe that the induced subgraph $G[Z]$ of $G$ on vertex set $Z$ is triangle-free.
 Also, $X$ and $Y$ are independent sets with no triangles contained in $G[X\cup Y]$.
@@
 Our program builds the clause $B$ and then checks its satisfiability.
-The proof of Theorem 12 follows because our program found a good coloring of $\textrm{KG}(8,2)$ (this coloring is given explicitly in Appendix A), but declared that the graphs $\textrm{KG}(9,2)$ and $\textrm{KG}(13,3)$ do not have good colorings.
-Our code is given in Appendix B.
+The proof of Theorem 12 for $R^{\textrm{KG}}_{2}(3,3)$ follows because our program found a good coloring of $\textrm{KG}(8,2)$, but declared that the graph $\textrm{KG}(9,2)$ does not have a good coloring.
+The bounds for the other Kneser Ramsey numbers in Theorem 12 were obtained in similar ways.
+The explicit colorings of $\textrm{KG}(8,2)$ with no monochromatic $K_{3}$, and of $\textrm{KG}(11,2)$ with no red $K_{4}$ or blue $K_{3}$ are given in Appendix A. Our code is given in Appendix B.
 4 Pálvölgyi’s problem
-We begin with a general discussion on monochromatic induced Kneser subgraphs, which will be useful in making progress on Problem 13.
-Let $r>1$ be an integer and let $\textrm{KG}(s,r)$ and $\textrm{KG}(t,r)$ be Kneser graphs.
-Define the induced $r$-Kneser Ramsey number $R^{\textrm{KG}}_{r,ind}(s,t)$ to be the minimum $n$ (if it exists) such that every red/blue edge-coloring of $\textrm{KG}(n,r)$ contains an induced red $\textrm{KG}(s,r)$ or an induced blue $\textrm{KG}(t,r)$.
-Theorem 18.
-For $r\geq 2$, $R^{\textrm{KG}}_{r,ind}(s,t)$ does not exists; that is, there exists a red/blue edge-coloring of $\textrm{KG}(n,r)$ with no red induced $\textrm{KG}(s,r)$ or blue induced $\textrm{KG}(t,r)$ .
-In the proof of Theorem 18 we use the set-system theorem of Bollobás.
+4.1 Induced Kneser graphs
+We start with a discussion on induced graphs of the Kneser graph.
+Recall that given integer $n\geq s\geq r$, we defined an induced $\textrm{KG}(s,r)$ as a subgraph $H$ of the graph $\textrm{KG}(n,r)$ induced on a vertex set $S\choose r$}, where $S\subset[n]$ is a ground set of size $s$.
+We first show that every induced graph $H$ of $\textrm{KG}(n,r)$ isomorphic to $\textrm{KG}(2r,r)$ is, in fact, an induced $\textrm{KG}(2r,r)$:
+Proposition 17.
+Suppose $H$ is an induced subgraph of $G=\textrm{KG}(n,r)$ isomorphic to $\textrm{KG}(2r,r)$.
+Then $H$ is an induced $\textrm{KG}(2r,r)$, that is, there exists a ground set $S\subset[n]$ of size $2r$ such that $H=G[{S\choose r}$].
+The proof will follow from the set-system theorem of Bollobás.
 Theorem 18 (Bollobás [2]).
@@
 Furthermore, if $m=\binom{a+b}{a}$ then there is some set $S$ of cardinality $a+b$ such that the sets $A_{i}$ are all the subsets of $S$ of size $a$, and $B_{i}=S\setminus A_{i}$ for all $i$.
-Proof of Theorem 18.
-We first show $R^{\textrm{KG}}_{r,ind}(2r,2r)$ does not exist. Observe that $\textrm{KG}(2r,r)$ is a matching of size $\frac{1}{2}\binom{2r}{r}$.
-Now, in any $\textrm{KG}(n,r)$ with $n\geq 2r$, we claim that every maximum induced matching $M$ has vertex set $\binom{S}{r}$, where $|S|=2r$.
-To this end, let $M$ be an induced matching with edge set $E=\{(X_{i},Y_{i})\colon 1\leq i\leq m\}$.
-Note that we must have $X_{i}\cap Y_{i}=\emptyset$ for all $i$, and $X_{i}\cap Y_{j}\neq\emptyset$ for all $i\neq j$ in order for $M$ to be an induced matching.
-Let $A_{i}=X_{i}$ for $1\leq i\leq m$, $A_{i}=Y_{i-m}$ for $m+1\leq i\leq 2m$, and let $B_{i}=Y_{i}$ for $1\leq i\leq m$, $B_{i}=X_{i-m}$ for $m+1\leq i\leq 2m$.
-Then the set pair system $(A_{i},B_{i})$, $1\leq i\leq 2m$ satisfies $A_{i}\cap B_{i}=\emptyset$ for all $i$ and $A_{i}\cap B_{j}\neq\emptyset$ for $i\neq j$.
-Thus, we may invoke Theorem 19, with $a=b=r$, and
-$2\cdot|E(M)|=2m\leq\binom{2r}{r}$.
-In fact, since we know there exists an induced matching of size $\frac{1}{2}\binom{2r}{r}$, equality can be achieved, and there is a set $S$ such that $|S|=2r$ where $A_{i}\subseteq S$ and $B_{i}=S\setminus A_{i}$ for each $i$.
-Hence, one edge in a maximum induced matching $M$ uniquely determines the $2r$ elements of the ground set which can be used to form $M$.
-Since a maximum induced matching $M$ can be uniquely determined by a single edge in $\textrm{KG}(n,r)$, this implies any two maximum induced matchings, i.e.
-matchings isomorphic to $\textrm{KG}(2r,r)$, are disjoint, and we can color each maximum induced matching bichromatically.
-Thus, we have a coloring of any $\textrm{KG}(n,r)$ with no monochromatic $\textrm{KG}(2r,r)$; i.e.
-$R^{\textrm{KG}}_{r,ind}(2r,2r)$ does not exist.
-Finally, observe that for any $s,t\geq 2r$, $\textrm{KG}(s,r)$ and $\textrm{KG}(t,r)$ each contain a subgraph isomorphic to $\textrm{KG}(2r,r)$.
-Hence, since there exists a red/blue edge-coloring of $\textrm{KG}(n,r)$ with no induced monochromatic $\textrm{KG}(2r,r)$ for any $n\geq 2r$, in the same edge-coloring there are no monochromatic induced $\textrm{KG}(s,r)$ nor monochromatic induced $\textrm{KG}(t,r)$.
+Proof of Proposition 17.
+Observe that $\textrm{KG}(2r,r)$ is a matching of size $\frac{1}{2}\binom{2r}{r}$.
+We claim that every induced matching in $\textrm{KG}(n,r)$ of size $\frac{1}{2}\binom{2r}{r}$ has vertex set $\binom{S}{r}$, for some $S\subset[n]$ and $|S|=2r$.
+Indeed, let $M=\{(X_{i},Y_{i})\colon 1\leq i\leq m\}$ be an induced matching of $\textrm{KG}(n,r)$.
+Then $X_{i}\cap Y_{i}=\emptyset$ for all $i\in[m]$, and $X_{i}\cap Y_{j}\neq\emptyset$ for all $i\neq j$.
+Set
+[…]
+and
+[…]
+Then the set pair system $\{(A_{i},B_{i})\mid 1\leq i\leq 2m\}$ satisfies the conditions of Theorem 18. Therefore, by the theorem,
+$2|M|=2m\leq\binom{2r}{r}$,
+and moreover, equality is achieved when the sets $A_{i}$ are all the $r$ subsets of the set $S=A_{1}\cup B_{1}$, where $B_{i}=S\setminus A_{i}$ for all $i$.
+This implies that $H$ is a induced $KG(2r,r)$ on ground set $S$, as claimed.
 ∎
-Proof of Theorem 14.
-Set $n=R_{r}(3r,3r)-1$.
-We aim to show that there is a red/blue coloring $c$ of $G=\textrm{KG}(n,r)$ containing no red induced $\textrm{KG}(3r,r)$ and no blue triangle.
-We begin by partitioning $V(G)$ into two sets.
-Consider a red/blue edge coloring $c^{\prime}$ of $G^{\prime}=K^{r}_{n}$ containing no monochromatic $K^{r}_{3r}$.
-Observe that $E(G^{\prime})=V(G)$ and there is a clear bijection between the two sets.
-Let us partition $V(G)$ into two sets $P_{1},P_{2}$ as follows: For $e=\{v_{1},v_{2},\dots,v_{r}\}\in E(G^{\prime})$, if $c(e)=R$, add $\{v_{1},v_{2},\dots,v_{r}\}$ to $P_{1}$.
-Otherwise, add $\{v_{1},v_{2},\dots,v_{r}\}$ to $P_{2}$.
-Observe that by the choice of $c^{\prime}$, neither $G[P_{1}]$ nor $G[P_{2}]$ contain an induced copy of $\textrm{KG}(3r,r)$ (since otherwise, $c^{\prime}$ would contain a monochromatic $K^{r}_{3r}$).
-Therefore, for any edge $e$ within $G[P_{1}]$ or $G[P_{2}]$, set $c(e)=R$ and otherwise, set $c(e)=B$.
+We return now to Pálvölgyi’s problem.
+In Problem 13 one would like to find a red/blue edge-coloring of $G=\textrm{KG}(n,r)$ with no red induced $\textrm{KG}(3r,r)$ and no blue triangle.
+We first show the existence of a red/blue edge-coloring of $G$ with no monochromatic induced Kneser graphs $\textrm{KG}(s,r)$ for all $n\geq s\geq 2r$.
+Proposition 19.
+Let $n\geq 2r\geq 4$ be integers. Then there exists a red/blue edge-coloring of $\textrm{KG}(n,r)$ with no monochromatic induced $\textrm{KG}(s,r)$ for all $n\geq s\geq 2r$.
+Proof.
+We say that an edge $XY\in\textrm{KG}(n,r)$ is interlacing, where $X,Y\in{[n]\choose r}$ are disjoint sets, if there exist $x,x^{\prime}\in X$ and $y,y^{\prime}\in Y$ such that either $y<x<y^{\prime}$ or $x<y<x^{\prime}$.
+Otherwise, we say that $XY$ is non-interlacing.
+Now color the interlacing edges of $\textrm{KG}(n,r)$ red, and the non-interlacing edges blue.
+We claim that in this coloring there is no monochromatic $\textrm{KG}(s,r)$ for any $n\geq s\geq 2r$.
+Indeed, suppose $H$ is an induced $\textrm{KG}(s,r)$ on ground set $S=\{z_{1}<z_{2}<\dots<z_{s}\}\subset[n]$ where $s\geq 2r\geq 4$.
+Then the edge $XY\in H$, where $X=\{z_{1},\dots z_{r}\}$ and $Y=\{z_{r+1},\dots z_{2r}\}$ is non-interlacing, while the edge $X^{\prime}Y^{\prime}\in H$, where $X^{\prime}=\{z_{2},\dots z_{r+1}\}$ and $Y^{\prime}=\{z_{1},z_{r+2}\dots z_{2r}\}$ is interlacing, showing that $H$ is not monochromatic.
+∎
+The proof shows that an induced $\textrm{KG}(s,r)$ is not monochromatic because it contains an induced $\textrm{KG}(2r,r)$, which is already not monochromatic under the edge coloring in the proof.
+Therefore, by Proposition 17 we can state Proposition 19 in a slightly stronger manner:
+Corollary 20.
+Let $n\geq 2r\geq 4$ be integers.
+There exists a red/blue edge-coloring of $G=\textrm{KG}(n,r)$ such that for all $n\geq s\geq 2r$, if $H$ is an induced subgraph of $G$ isomorphic to $\textrm{KG}(s,r)$, then $H$ is not monochromatic.
+Note that Problem 13 is not resolved by Corollary 20 because a triangle is an induced Kneser graph only if $r=1$. In the next subsection we prove Theorem 14, which gives a lower bound for Problem 13.
+4.2 Proof of Theorem 14
+Set $n=R_{r}(3r,3r)-1$, where $R_{r}(3r,3r)$ is the minimum integer $m$ such that in any red/blue coloring of the edges of the complete $r$-uniform hypergraph on $m$ vertices $K^{r}_{m}$ there is a monochromatic $K^{r}_{3r}$.
+We would like to show that there is a red/blue coloring $c$ of $G=\textrm{KG}(n,r)$ containing no red induced $\textrm{KG}(3r,r)$ and no blue triangle.
+Let $c^{\prime}$ be a red/blue edge coloring of the graph $G^{\prime}=K^{r}_{n}$ containing no monochromatic $K^{r}_{3r}$ ($c^{\prime}$ exists by the choice of $n$).
+We partition $E(G^{\prime})=V(G)$ into two sets $P_{1},P_{2}$ as follows: for $e\in E(G^{\prime})$, if $c^{\prime}(e)=R$, then $e\in P_{1}$.
+Otherwise, $e\in P_{2}$.
+Observe that by the choice of $c^{\prime}$, neither $G[P_{1}]$ nor $G[P_{2}]$ contain an induced copy of $\textrm{KG}(3r,r)$.
+Indeed, if $G[P_{i}]$ contains an induced copy of $\textrm{KG}(3r,r)$, then there exists $S\subset[n]$ of size $3r$ such that ${S\choose r}\subset P_{i}$.
+But this implies that $c^{\prime}$ contains a monochromatic copy of $K^{r}_{3r}$ on vertex set $S$.
+We now define $c:E(G)\to\{R,B\}$ as follows.
+For any edge $e\in G[P_{R}]$ or $e\in G[P_{B}]$, set $c(e)=R$.
+Otherwise, set $c(e)=B$.
 Since the blue subgraph is a bipartite graph, it is triangle-free.
-And so, $c$ is a red/blue edge coloring of $G$ containing no red induced $\textrm{KG}(3r,r)$ and no blue triangle.
+Hence $c$ is a red/blue edge coloring of $G$ containing no red induced $\textrm{KG}(3r,r)$ and no blue triangle.
 ∎
@@
 $R^{\textrm{KG}}_{r}(s_{1},\dots,s_{t})\leq(m_{1}+1)r+\sum_{i=2}^{t}(m_{i}-1).$
-Appendix A: A good coloring of $\textrm{KG}(8,2)$.
-Here, we provide a coloring of $\textrm{KG}(8,2)$ with no monochromatic triangle.
-The red edges are:
-((1,2),(3,4)),((1,2),(3,5)),((1,2),(3,6)),((1,2),(3,7)),((1,2),(3,8)),((1,3),(2,4)),((1,3),(2,5)),((1,3),(2,6)),((1,3),(2,7)),((1,3),(2,8)),((1,4),(2,3)),((1,4),(2,6)),((1,4),(2,8)),((1,4),(3,6)),((1,4),(3,8)),((1,4),(5,6)),((1,4),(6,8)),((1,5),(2,3)),((1,5),(2,6)),((1,5),(2,7)),((1,5),(3,6)),((1,5),(3,7)),((1,5),(6,7)),((1,5),(6,8)),((1,5),(7,8)),((1,6),(2,3)),((1,6),(2,4)),((1,6),(2,5)),((1,6),(3,4)),((1,6),(3,5)),((1,6),(4,5)),((1,6),(5,7)),((1,6),(5,8)),((1,7),(2,3)),((1,7),(2,5)),((1,7),(2,8)),((1,7),(3,5)),((1,7),(3,8)),((1,7),(4,8)),((1,7),(5,8)),((1,8),(2,3)),((1,8),(2,4)),((1,8),(2,7)),((1,8),(3,4)),((1,8),(3,7)),((1,8),(4,6)),((1,8),(4,7)),((1,8),(6,7)),((2,4),(3,6)),((2,4),(3,8)),((2,4),(5,6)),((2,4),(6,8)),((2,5),(3,6)),((2,5),(3,7)),((2,5),(6,7)),((2,5),(6,8)),((2,5),(7,8)),((2,6),(3,4)),((2,6),(3,5)),((2,6),(4,5)),((2,6),(5,7)),((2,6),(5,8)),((2,7),(3,5)),((2,7),(3,8)),((2,7),(4,5)),((2,7),(4,8)),((2,7),(5,8)),((2,8),(3,4)),((2,8),(3,7)),((2,8),(4,5)),((2,8),(4,6)),((2,8),(4,7)),((3,4),(5,6)),((3,4),(6,8)),((3,5),(6,7)),((3,5),(6,8)),((3,5),(7,8)),((3,6),(4,5)),((3,6),(5,7)),((3,6),(5,8)),((3,7),(4,8)),((3,7),(5,8)),((3,8),(4,6)),((3,8),(4,7)),((3,8),(6,7)),((4,5),(6,7)),((4,5),(6,8)),((4,5),(7,8)),((4,6),(5,7)),((4,6),(5,8)),((4,6),(7,8)),((4,7),(5,6)),((4,7),(5,8)),((4,7),(6,8)),((4,8),(5,6)),((4,8),(5,7)),((4,8),(6,7)),((5,6),(7,8)),((5,7),(6,8)),((5,8),(6,7))((1,2),(3,4)),((1,2),(3,5)),((1,2),(3,6)),((1,2),(3,7)),((1,2),(3,8)),((1,3),(2,4)),\\ ((1,3),(2,5)),((1,3),(2,6)),((1,3),(2,7)),((1,3),(2,8)),((1,4),(2,3)),((1,4),(2,6)),\\ ((1,4),(2,8)),((1,4),(3,6)),((1,4),(3,8)),((1,4),(5,6)),((1,4),(6,8)),((1,5),(2,3)),\\ ((1,5),(2,6)),((1,5),(2,7)),((1,5),(3,6)),((1,5),(3,7)),((1,5),(6,7)),((1,5),(6,8)),\\ ((1,5),(7,8)),((1,6),(2,3)),((1,6),(2,4)),((1,6),(2,5)),((1,6),(3,4)),((1,6),(3,5)),\\ ((1,6),(4,5)),((1,6),(5,7)),((1,6),(5,8)),((1,7),(2,3)),((1,7),(2,5)),((1,7),(2,8)),\\ ((1,7),(3,5)),((1,7),(3,8)),((1,7),(4,8)),((1,7),(5,8)),((1,8),(2,3)),((1,8),(2,4)),\\ ((1,8),(2,7)),((1,8),(3,4)),((1,8),(3,7)),((1,8),(4,6)),((1,8),(4,7)),((1,8),(6,7)),\\ ((2,4),(3,6)),((2,4),(3,8)),((2,4),(5,6)),((2,4),(6,8)),((2,5),(3,6)),((2,5),(3,7)),\\ ((2,5),(6,7)),((2,5),(6,8)),((2,5),(7,8)),((2,6),(3,4)),((2,6),(3,5)),((2,6),(4,5)),\\ ((2,6),(5,7)),((2,6),(5,8)),((2,7),(3,5)),((2,7),(3,8)),((2,7),(4,5)),((2,7),(4,8)),\\ ((2,7),(5,8)),((2,8),(3,4)),((2,8),(3,7)),((2,8),(4,5)),((2,8),(4,6)),((2,8),(4,7)),\\ ((3,4),(5,6)),((3,4),(6,8)),((3,5),(6,7)),((3,5),(6,8)),((3,5),(7,8)),((3,6),(4,5)),\\ ((3,6),(5,7)),((3,6),(5,8)),((3,7),(4,8)),((3,7),(5,8)),((3,8),(4,6)),((3,8),(4,7)),\\ ((3,8),(6,7)),((4,5),(6,7)),((4,5),(6,8)),((4,5),(7,8)),((4,6),(5,7)),((4,6),(5,8)),\\ ((4,6),(7,8)),((4,7),(5,6)),((4,7),(5,8)),((4,7),(6,8)),((4,8),(5,6)),((4,8),(5,7)),\\ ((4,8),(6,7)),((5,6),(7,8)),((5,7),(6,8)),((5,8),(6,7))
-The remaining edges are blue.
+Appendix A Appendix: Explicit colorings.
+A.1 A good coloring of $\textrm{KG}(8,2)$
+Here we provide a coloring of $\textrm{KG}(8,2)$ with no monochromatic $K_{3}$. This shows $R^{\textrm{KG}}_{2}(3,3)\geq 9$.
 We can describe part of this coloring as follows.
@@
 Then all the edges from $A$ to $C$ as well as all the edges within $B$ are red, and all the edges from $A$ to $B$ are blue.
-Appendix B: Our code
+The red edges are:
+Edges from $A$ to $C$:
+((1,2),(3,4)),((1,2),(3,5)),((1,2),(3,6)),((1,2),(3,7)),((1,2),(3,8)),((1,3),(2,4)),((1,3),(2,5)),((1,3),(2,6)),((1,3),(2,7)),((1,3),(2,8)),((1,4),(2,3)),((1,5),(2,3)),((1,6),(2,3)),((1,7),(2,3)),((1,8),(2,3))((1,2),(3,4)),((1,2),(3,5)),((1,2),(3,6)),((1,2),(3,7)),((1,2),(3,8)),((1,3),(2,4)),\\ ((1,3),(2,5)),((1,3),(2,6)),((1,3),(2,7)),((1,3),(2,8)),((1,4),(2,3)),((1,5),(2,3)),\\ ((1,6),(2,3)),((1,7),(2,3)),((1,8),(2,3))
+Edges from $B$ to $C$:
+((1,4),(5,6)),((1,4),(6,8)),((1,5),(6,7)),((1,5),(6,8)),((1,5),(7,8)),((1,6),(4,5)),((1,6),(5,7)),((1,6),(5,8)),((1,7),(4,8)),((1,7),(5,8)),((1,8),(4,6)),((1,8),(4,7)),((1,8),(6,7)),((2,4),(5,6)),((2,4),(6,8)),((2,5),(6,7)),((2,5),(6,8)),((2,5),(7,8)),((2,6),(4,5)),((2,6),(5,7)),((2,6),(5,8)),((2,7),(4,5)),((2,7),(4,8)),((2,7),(5,8)),((2,8),(4,5)),((2,8),(4,6)),((2,8),(4,7)),((3,4),(5,6)),((3,4),(6,8)),((3,5),(6,7)),((3,5),(6,8)),((3,5),(7,8)),((3,6),(4,5)),((3,6),(5,7)),((3,6),(5,8)),((3,7),(4,8)),((3,7),(5,8)),((3,8),(4,6)),((3,8),(4,7)),((3,8),(6,7))((1,4),(5,6)),((1,4),(6,8)),((1,5),(6,7)),((1,5),(6,8)),((1,5),(7,8)),((1,6),(4,5)),\\ ((1,6),(5,7)),((1,6),(5,8)),((1,7),(4,8)),((1,7),(5,8)),((1,8),(4,6)),((1,8),(4,7)),\\ ((1,8),(6,7)),((2,4),(5,6)),((2,4),(6,8)),((2,5),(6,7)),((2,5),(6,8)),((2,5),(7,8)),\\ ((2,6),(4,5)),((2,6),(5,7)),((2,6),(5,8)),((2,7),(4,5)),((2,7),(4,8)),((2,7),(5,8)),\\ ((2,8),(4,5)),((2,8),(4,6)),((2,8),(4,7)),((3,4),(5,6)),((3,4),(6,8)),((3,5),(6,7)),\\ ((3,5),(6,8)),((3,5),(7,8)),((3,6),(4,5)),((3,6),(5,7)),((3,6),(5,8)),((3,7),(4,8)),\\ ((3,7),(5,8)),((3,8),(4,6)),((3,8),(4,7)),((3,8),(6,7))
+Edges within $B$:
+((4,5),(6,7)),((4,5),(6,8)),((4,5),(7,8)),((4,6),(5,7)),((4,6),(5,8)),((4,6),(7,8)),((4,7),(5,6)),((4,7),(5,8)),((4,7),(6,8)),((4,8),(5,6)),((4,8),(5,7)),((4,8),(6,7)),((5,6),(7,8)),((5,7),(6,8)),((5,8),(6,7))((4,5),(6,7)),((4,5),(6,8)),((4,5),(7,8)),((4,6),(5,7)),((4,6),(5,8)),((4,6),(7,8)),\\ ((4,7),(5,6)),((4,7),(5,8)),((4,7),(6,8)),((4,8),(5,6)),((4,8),(5,7)),((4,8),(6,7)),\\ ((5,6),(7,8)),((5,7),(6,8)),((5,8),(6,7))
+Edges within $C$:
+((1,4),(2,6)),((1,4),(2,8)),((1,4),(3,6)),((1,4),(3,8)),((1,5),(2,6)),((1,5),(2,7)),((1,5),(3,6)),((1,5),(3,7)),((1,6),(2,4)),((1,6),(2,5)),((1,6),(3,4)),((1,6),(3,5)),((1,7),(2,5)),((1,7),(2,8)),((1,7),(3,5)),((1,7),(3,8)),((1,8),(2,4)),((1,8),(2,7)),((1,8),(3,4)),((1,8),(3,7)),((2,4),(3,6)),((2,4),(3,8)),((2,5),(3,6)),((2,5),(3,7)),((2,6),(3,4)),((2,6),(3,5)),((2,7),(3,5)),((2,7),(3,8)),((2,8),(3,4)),((2,8),(3,7))((1,4),(2,6)),((1,4),(2,8)),((1,4),(3,6)),((1,4),(3,8)),((1,5),(2,6)),((1,5),(2,7)),\\ ((1,5),(3,6)),((1,5),(3,7)),((1,6),(2,4)),((1,6),(2,5)),((1,6),(3,4)),((1,6),(3,5)),\\ ((1,7),(2,5)),((1,7),(2,8)),((1,7),(3,5)),((1,7),(3,8)),((1,8),(2,4)),((1,8),(2,7)),\\ ((1,8),(3,4)),((1,8),(3,7)),((2,4),(3,6)),((2,4),(3,8)),((2,5),(3,6)),((2,5),(3,7)),\\ ((2,6),(3,4)),((2,6),(3,5)),((2,7),(3,5)),((2,7),(3,8)),((2,8),(3,4)),((2,8),(3,7))
+The remaining edges are blue.
+A.2 A good coloring of $\textrm{KG}(11,2)$
+Below is a coloring of $\textrm{KG}(11,2)$ with no red $K_{4}$ or blue $K_{3}$. This shows $R^{\textrm{KG}}_{2}(4,3)\geq 12$.
+The red edges are:
+((1,2),(3,5)),((1,2),(3,9)),((1,2),(3,10)),((1,2),(4,5)),((1,2),(4,10)),((1,2),(5,6)),((1,2),(5,7))((1,2),(5,8)),((1,2),(5,9)),((1,2),(5,10)),((1,2),(5,11)),((1,2),(6,10)),((1,2),(8,10)),((1,2),(9,10))((1,2),(10,11)),((1,3),(2,7)),((1,3),(2,9)),((1,3),(4,9)),((1,3),(5,9)),((1,3),(6,7)),((1,3),(6,9))((1,3),(7,8)),((1,3),(7,9)),((1,3),(7,11)),((1,3),(8,9)),((1,3),(9,11)),((1,4),(2,5)),((1,4),(2,8))((1,4),(2,11)),((1,4),(3,9)),((1,4),(3,11)),((1,4),(5,6)),((1,4),(5,8)),((1,4),(5,9)),((1,4),(5,11))((1,4),(6,8)),((1,4),(6,11)),((1,4),(8,9)),((1,4),(8,11)),((1,4),(9,11)),((1,5),(2,3)),((1,5),(2,4))((1,5),(2,6)),((1,5),(2,8)),((1,5),(2,10)),((1,5),(2,11)),((1,5),(3,6)),((1,5),(3,8)),((1,5),(3,9))((1,5),(3,11)),((1,5),(4,6)),((1,5),(4,8)),((1,5),(4,11)),((1,6),(2,3)),((1,6),(2,5)),((1,6),(2,10))((1,6),(3,5)),((1,6),(3,7)),((1,6),(3,9)),((1,6),(4,5)),((1,6),(4,10)),((1,6),(5,8)),((1,6),(5,9))((1,6),(5,10)),((1,6),(5,11)),((1,6),(8,10)),((1,6),(9,10)),((1,6),(10,11)),((1,7),(2,3)),((1,7),(2,4))((1,7),(2,6)),((1,7),(2,8)),((1,7),(2,11)),((1,7),(3,4)),((1,7),(3,6)),((1,7),(3,8)),((1,7),(3,9))((1,7),(3,11)),((1,7),(4,6)),((1,7),(4,8)),((1,7),(4,11)),((1,7),(6,8)),((1,7),(6,11)),((1,7),(8,11))((1,8),(2,5)),((1,8),(2,10)),((1,8),(3,5)),((1,8),(3,6)),((1,8),(3,9)),((1,8),(4,5)),((1,8),(5,6))((1,8),(5,9)),((1,8),(5,10)),((1,8),(5,11)),((1,8),(6,10)),((1,8),(9,10)),((1,8),(10,11)),((1,9),(2,3))((1,9),(2,10)),((1,9),(3,4)),((1,9),(3,5)),((1,9),(3,6)),((1,9),(3,7)),((1,9),(3,8)),((1,9),(3,10))((1,9),(3,11)),((1,9),(4,10)),((1,9),(5,8)),((1,9),(5,11)),((1,9),(6,10)),((1,9),(8,10)),((1,9),(10,11))((1,10),(2,3)),((1,10),(2,5)),((1,10),(2,6)),((1,10),(2,8)),((1,10),(2,9)),((1,10),(2,11)),((1,10),(3,5))((1,10),(3,6)),((1,10),(3,8)),((1,10),(3,9)),((1,10),(5,6)),((1,10),(5,8)),((1,10),(5,9)),((1,10),(5,11))((1,10),(6,8)),((1,10),(6,9)),((1,11),(2,5)),((1,11),(2,10)),((1,11),(3,5)),((1,11),(3,9)),((1,11),(4,5))((1,11),(5,6)),((1,11),(5,7)),((1,11),(5,8)),((1,11),(5,9)),((1,11),(5,10)),((1,11),(6,10)),((1,11),(8,10))((1,11),(9,10)),((2,3),(4,7)),((2,3),(4,9)),((2,3),(5,7)),((2,3),(6,7)),((2,3),(7,8)),((2,3),(7,9))((2,3),(7,10)),((2,3),(7,11)),((2,3),(8,9)),((2,4),(5,6)),((2,4),(5,7)),((2,4),(5,8)),((2,4),(5,9))((2,4),(5,11)),((2,4),(6,7)),((2,4),(6,9)),((2,4),(7,8)),((2,4),(7,9)),((2,4),(7,11)),((2,4),(8,9))((2,4),(9,11)),((2,5),(3,4)),((2,5),(4,6)),((2,5),(4,7)),((2,5),(4,8)),((2,5),(4,9)),((2,5),(4,11))((2,5),(7,10)),((2,6),(3,7)),((2,6),(3,10)),((2,6),(4,7)),((2,6),(4,10)),((2,6),(5,7)),((2,6),(5,10))((2,6),(7,8)),((2,6),(7,9)),((2,6),(7,10)),((2,6),(7,11)),((2,6),(8,10)),((2,6),(9,10)),((2,6),(10,11))((2,7),(3,4)),((2,7),(3,5)),((2,7),(3,6)),((2,7),(3,8)),((2,7),(3,9)),((2,7),(3,11)),((2,7),(4,6))((2,7),(4,8)),((2,7),(4,11)),((2,7),(6,8)),((2,7),(6,11)),((2,7),(8,11)),((2,8),(3,10)),((2,8),(4,5))((2,8),(4,7)),((2,8),(4,10)),((2,8),(5,7)),((2,8),(5,10)),((2,8),(6,7)),((2,8),(6,10)),((2,8),(7,10))((2,8),(7,11)),((2,8),(9,10)),((2,8),(10,11)),((2,9),(3,4)),((2,9),(3,10)),((2,9),(4,6)),((2,9),(4,7))((2,9),(4,8)),((2,9),(4,10)),((2,9),(4,11)),((2,9),(5,10)),((2,9),(6,10)),((2,9),(7,10)),((2,9),(8,10))((2,9),(10,11)),((2,10),(4,9)),((2,10),(6,8)),((2,10),(6,9)),((2,10),(6,11)),((2,10),(8,9)),((2,10),(8,11))((2,10),(9,11)),((2,11),(3,10)),((2,11),(4,7)),((2,11),(4,10)),((2,11),(5,7)),((2,11),(5,10)),((2,11),(6,7))((2,11),(6,10)),((2,11),(7,8)),((2,11),(7,9)),((2,11),(7,10)),((2,11),(8,10)),((2,11),(9,10)),((3,4),(5,7))((3,4),(5,9)),((3,4),(6,7)),((3,4),(6,9)),((3,4),(7,8)),((3,4),(7,9)),((3,4),(7,11)),((3,4),(8,9))((3,4),(9,11)),((3,5),(4,7)),((3,5),(4,9)),((3,5),(6,7)),((3,5),(7,8)),((3,5),(7,9)),((3,5),(7,10))((3,5),(7,11)),((3,6),(4,7)),((3,6),(4,9)),((3,6),(5,7)),((3,6),(7,8)),((3,6),(7,9)),((3,6),(7,10))((3,6),(7,11)),((3,7),(4,6)),((3,7),(4,9)),((3,7),(5,6)),((3,7),(6,8)),((3,7),(6,9)),((3,7),(6,11))((3,7),(8,11)),((3,8),(4,5)),((3,8),(4,7)),((3,8),(4,9)),((3,8),(5,7)),((3,8),(6,7)),((3,8),(7,9))((3,8),(7,10)),((3,8),(7,11)),((3,9),(6,7)),((3,9),(7,8)),((3,9),(7,11)),((3,10),(4,9)),((3,10),(6,8))((3,10),(6,9)),((3,10),(6,11)),((3,10),(8,9)),((3,10),(8,11)),((3,10),(9,11)),((3,11),(4,5)),((3,11),(4,7))((3,11),(4,9)),((3,11),(5,7)),((3,11),(6,7)),((3,11),(6,9)),((3,11),(7,8)),((3,11),(7,9)),((3,11),(7,10))((4,5),(6,8)),((4,5),(8,9)),((4,5),(8,11)),((4,5),(9,11)),((4,6),(5,7)),((4,6),(5,8)),((4,6),(5,9))((4,6),(5,11)),((4,6),(7,8)),((4,6),(7,9)),((4,6),(7,11)),((4,6),(8,9)),((4,6),(9,11)),((4,7),(5,6))((4,7),(5,8)),((4,7),(5,9)),((4,7),(5,11)),((4,7),(6,11)),((4,7),(8,11)),((4,7),(9,11)),((4,8),(5,6))((4,8),(5,7)),((4,8),(5,9)),((4,8),(5,11)),((4,8),(6,7)),((4,8),(6,9)),((4,8),(7,9)),((4,8),(7,11))((4,8),(9,11)),((4,9),(5,8)),((4,9),(5,11)),((4,9),(6,10)),((4,9),(8,10)),((4,9),(10,11)),((4,10),(6,8))((4,10),(6,9)),((4,10),(6,11)),((4,10),(8,9)),((4,10),(8,11)),((4,10),(9,11)),((4,11),(5,6)),((4,11),(5,7))((4,11),(5,8)),((4,11),(5,9)),((4,11),(6,7)),((4,11),(6,9)),((4,11),(7,8)),((4,11),(7,9)),((4,11),(8,9))((5,7),(6,11)),((5,7),(8,11)),((5,10),(6,8)),((5,10),(6,9)),((5,10),(6,11)),((5,10),(8,9)),((5,10),(8,11))((5,10),(9,11)),((6,7),(8,11)),((6,8),(7,10)),((6,8),(7,11)),((6,8),(9,10)),((6,8),(10,11)),((6,9),(8,10))((6,9),(10,11)),((6,10),(8,9)),((6,10),(8,11)),((6,10),(9,11)),((6,11),(7,8)),((6,11),(7,9)),((6,11),(7,10))((6,11),(8,10)),((6,11),(9,10)),((7,9),(8,11)),((7,10),(8,11)),((8,9),(10,11)),((8,10),(9,11)),((8,11),(9,10))((1,2),(3,5)),((1,2),(3,9)),((1,2),(3,10)),((1,2),(4,5)),((1,2),(4,10)),((1,2),(5,6)),((1,2),(5,7))\\ ((1,2),(5,8)),((1,2),(5,9)),((1,2),(5,10)),((1,2),(5,11)),((1,2),(6,10)),((1,2),(8,10)),((1,2),(9,10))\\ ((1,2),(10,11)),((1,3),(2,7)),((1,3),(2,9)),((1,3),(4,9)),((1,3),(5,9)),((1,3),(6,7)),((1,3),(6,9))\\ ((1,3),(7,8)),((1,3),(7,9)),((1,3),(7,11)),((1,3),(8,9)),((1,3),(9,11)),((1,4),(2,5)),((1,4),(2,8))\\ ((1,4),(2,11)),((1,4),(3,9)),((1,4),(3,11)),((1,4),(5,6)),((1,4),(5,8)),((1,4),(5,9)),((1,4),(5,11))\\ ((1,4),(6,8)),((1,4),(6,11)),((1,4),(8,9)),((1,4),(8,11)),((1,4),(9,11)),((1,5),(2,3)),((1,5),(2,4))\\ ((1,5),(2,6)),((1,5),(2,8)),((1,5),(2,10)),((1,5),(2,11)),((1,5),(3,6)),((1,5),(3,8)),((1,5),(3,9))\\ ((1,5),(3,11)),((1,5),(4,6)),((1,5),(4,8)),((1,5),(4,11)),((1,6),(2,3)),((1,6),(2,5)),((1,6),(2,10))\\ ((1,6),(3,5)),((1,6),(3,7)),((1,6),(3,9)),((1,6),(4,5)),((1,6),(4,10)),((1,6),(5,8)),((1,6),(5,9))\\ ((1,6),(5,10)),((1,6),(5,11)),((1,6),(8,10)),((1,6),(9,10)),((1,6),(10,11)),((1,7),(2,3)),((1,7),(2,4))\\ ((1,7),(2,6)),((1,7),(2,8)),((1,7),(2,11)),((1,7),(3,4)),((1,7),(3,6)),((1,7),(3,8)),((1,7),(3,9))\\ ((1,7),(3,11)),((1,7),(4,6)),((1,7),(4,8)),((1,7),(4,11)),((1,7),(6,8)),((1,7),(6,11)),((1,7),(8,11))\\ ((1,8),(2,5)),((1,8),(2,10)),((1,8),(3,5)),((1,8),(3,6)),((1,8),(3,9)),((1,8),(4,5)),((1,8),(5,6))\\ ((1,8),(5,9)),((1,8),(5,10)),((1,8),(5,11)),((1,8),(6,10)),((1,8),(9,10)),((1,8),(10,11)),((1,9),(2,3))\\ ((1,9),(2,10)),((1,9),(3,4)),((1,9),(3,5)),((1,9),(3,6)),((1,9),(3,7)),((1,9),(3,8)),((1,9),(3,10))\\ ((1,9),(3,11)),((1,9),(4,10)),((1,9),(5,8)),((1,9),(5,11)),((1,9),(6,10)),((1,9),(8,10)),((1,9),(10,11))\\ ((1,10),(2,3)),((1,10),(2,5)),((1,10),(2,6)),((1,10),(2,8)),((1,10),(2,9)),((1,10),(2,11)),((1,10),(3,5))\\ ((1,10),(3,6)),((1,10),(3,8)),((1,10),(3,9)),((1,10),(5,6)),((1,10),(5,8)),((1,10),(5,9)),((1,10),(5,11))\\ ((1,10),(6,8)),((1,10),(6,9)),((1,11),(2,5)),((1,11),(2,10)),((1,11),(3,5)),((1,11),(3,9)),((1,11),(4,5))\\ ((1,11),(5,6)),((1,11),(5,7)),((1,11),(5,8)),((1,11),(5,9)),((1,11),(5,10)),((1,11),(6,10)),((1,11),(8,10))\\ ((1,11),(9,10)),((2,3),(4,7)),((2,3),(4,9)),((2,3),(5,7)),((2,3),(6,7)),((2,3),(7,8)),((2,3),(7,9))\\ ((2,3),(7,10)),((2,3),(7,11)),((2,3),(8,9)),((2,4),(5,6)),((2,4),(5,7)),((2,4),(5,8)),((2,4),(5,9))\\ ((2,4),(5,11)),((2,4),(6,7)),((2,4),(6,9)),((2,4),(7,8)),((2,4),(7,9)),((2,4),(7,11)),((2,4),(8,9))\\ ((2,4),(9,11)),((2,5),(3,4)),((2,5),(4,6)),((2,5),(4,7)),((2,5),(4,8)),((2,5),(4,9)),((2,5),(4,11))\\ ((2,5),(7,10)),((2,6),(3,7)),((2,6),(3,10)),((2,6),(4,7)),((2,6),(4,10)),((2,6),(5,7)),((2,6),(5,10))\\ ((2,6),(7,8)),((2,6),(7,9)),((2,6),(7,10)),((2,6),(7,11)),((2,6),(8,10)),((2,6),(9,10)),((2,6),(10,11))\\ ((2,7),(3,4)),((2,7),(3,5)),((2,7),(3,6)),((2,7),(3,8)),((2,7),(3,9)),((2,7),(3,11)),((2,7),(4,6))\\ ((2,7),(4,8)),((2,7),(4,11)),((2,7),(6,8)),((2,7),(6,11)),((2,7),(8,11)),((2,8),(3,10)),((2,8),(4,5))\\ ((2,8),(4,7)),((2,8),(4,10)),((2,8),(5,7)),((2,8),(5,10)),((2,8),(6,7)),((2,8),(6,10)),((2,8),(7,10))\\ ((2,8),(7,11)),((2,8),(9,10)),((2,8),(10,11)),((2,9),(3,4)),((2,9),(3,10)),((2,9),(4,6)),((2,9),(4,7))\\ ((2,9),(4,8)),((2,9),(4,10)),((2,9),(4,11)),((2,9),(5,10)),((2,9),(6,10)),((2,9),(7,10)),((2,9),(8,10))\\ ((2,9),(10,11)),((2,10),(4,9)),((2,10),(6,8)),((2,10),(6,9)),((2,10),(6,11)),((2,10),(8,9)),((2,10),(8,11))\\ ((2,10),(9,11)),((2,11),(3,10)),((2,11),(4,7)),((2,11),(4,10)),((2,11),(5,7)),((2,11),(5,10)),((2,11),(6,7))\\ ((2,11),(6,10)),((2,11),(7,8)),((2,11),(7,9)),((2,11),(7,10)),((2,11),(8,10)),((2,11),(9,10)),((3,4),(5,7))\\ ((3,4),(5,9)),((3,4),(6,7)),((3,4),(6,9)),((3,4),(7,8)),((3,4),(7,9)),((3,4),(7,11)),((3,4),(8,9))\\ ((3,4),(9,11)),((3,5),(4,7)),((3,5),(4,9)),((3,5),(6,7)),((3,5),(7,8)),((3,5),(7,9)),((3,5),(7,10))\\ ((3,5),(7,11)),((3,6),(4,7)),((3,6),(4,9)),((3,6),(5,7)),((3,6),(7,8)),((3,6),(7,9)),((3,6),(7,10))\\ ((3,6),(7,11)),((3,7),(4,6)),((3,7),(4,9)),((3,7),(5,6)),((3,7),(6,8)),((3,7),(6,9)),((3,7),(6,11))\\ ((3,7),(8,11)),((3,8),(4,5)),((3,8),(4,7)),((3,8),(4,9)),((3,8),(5,7)),((3,8),(6,7)),((3,8),(7,9))\\ ((3,8),(7,10)),((3,8),(7,11)),((3,9),(6,7)),((3,9),(7,8)),((3,9),(7,11)),((3,10),(4,9)),((3,10),(6,8))\\ ((3,10),(6,9)),((3,10),(6,11)),((3,10),(8,9)),((3,10),(8,11)),((3,10),(9,11)),((3,11),(4,5)),((3,11),(4,7))\\ ((3,11),(4,9)),((3,11),(5,7)),((3,11),(6,7)),((3,11),(6,9)),((3,11),(7,8)),((3,11),(7,9)),((3,11),(7,10))\\ ((4,5),(6,8)),((4,5),(8,9)),((4,5),(8,11)),((4,5),(9,11)),((4,6),(5,7)),((4,6),(5,8)),((4,6),(5,9))\\ ((4,6),(5,11)),((4,6),(7,8)),((4,6),(7,9)),((4,6),(7,11)),((4,6),(8,9)),((4,6),(9,11)),((4,7),(5,6))\\ ((4,7),(5,8)),((4,7),(5,9)),((4,7),(5,11)),((4,7),(6,11)),((4,7),(8,11)),((4,7),(9,11)),((4,8),(5,6))\\ ((4,8),(5,7)),((4,8),(5,9)),((4,8),(5,11)),((4,8),(6,7)),((4,8),(6,9)),((4,8),(7,9)),((4,8),(7,11))\\ ((4,8),(9,11)),((4,9),(5,8)),((4,9),(5,11)),((4,9),(6,10)),((4,9),(8,10)),((4,9),(10,11)),((4,10),(6,8))\\ ((4,10),(6,9)),((4,10),(6,11)),((4,10),(8,9)),((4,10),(8,11)),((4,10),(9,11)),((4,11),(5,6)),((4,11),(5,7))\\ ((4,11),(5,8)),((4,11),(5,9)),((4,11),(6,7)),((4,11),(6,9)),((4,11),(7,8)),((4,11),(7,9)),((4,11),(8,9))\\ ((5,7),(6,11)),((5,7),(8,11)),((5,10),(6,8)),((5,10),(6,9)),((5,10),(6,11)),((5,10),(8,9)),((5,10),(8,11))\\ ((5,10),(9,11)),((6,7),(8,11)),((6,8),(7,10)),((6,8),(7,11)),((6,8),(9,10)),((6,8),(10,11)),((6,9),(8,10))\\ ((6,9),(10,11)),((6,10),(8,9)),((6,10),(8,11)),((6,10),(9,11)),((6,11),(7,8)),((6,11),(7,9)),((6,11),(7,10))\\ ((6,11),(8,10)),((6,11),(9,10)),((7,9),(8,11)),((7,10),(8,11)),((8,9),(10,11)),((8,10),(9,11)),((8,11),(9,10))
+The rest of the edges are blue.
+Appendix B Appendix: Our code
 Below is the code we used to obtain the results in Theorem 12.
     import itertools as it
-    from pysat.solvers import Glucose3
+    from pysat.solvers import Glucose42
     Builds Kneser Graph Kn(n, r)
@@
     For no red K_s, for every s-clique, need at least 1 blue edge.
-    That is, if e_1, ..., e_k are the edges of an s-clique and e_i = 1 if red, e_i = -1 if blue,
-    want to satisfy (-e_1 v -e_2 v ... v -e_k)
+    That is, if e_1, ..., e_k are the edges of an s-clique and e_i = 1 if red,
+    e_i = -1 if blue, want to satisfy (-e_1 v -e_2 v ... v -e_k)
     For no blue K_t, for every t-clique, need at least 1 red edge.
-    That is, if e_1, ..., e_k are the edges of a t-clique and e_i = 1 if red, e_i = -1 if blue,
-    want to satisfy (-e_1 v -e_2 v ... v -e_k)
+    That is, if e_1, ..., e_k are the edges of a t-clique and e_i = 1 if red,
+    e_i = -1 if blue, want to satisfy (-e_1 v -e_2 v ... v -e_k)
     (NOTE: 1 corresponds to True, -1 corresponds to False)
-    If there is a good coloring, outputs (True, R, B), where R = red edges, B = blue edges
+    If there is a good coloring, outputs (True, R, B), where R = red edges,
+    B = blue edges
     If there is no good coloring, outputs (False, [], [])
     def build_sat_instances(n, r, s, t):
-        solver = Glucose3()
+        solver = Glucose42()
         clauses = []
@@
             #check that the s-tuple forms a clique in Kn(n,r)
-            if all([set(s_tuple[i]).isdisjoint(set(s_tuple[j])) for (i,j) in it.combinations(range(len(s_tuple)), 2)]):
+            if all([set(s_tuple[i]).isdisjoint(set(s_tuple[j])) for (i,j) in
+            it.combinations(range(len(s_tuple)), 2)]):
                 #Add clause forbidding monochromatic K_s and K_t (if s = t)
@@
                 #check that the t-tuple forms a clique in Kn(n,r)
-                if all([set(t_tuple[i]).isdisjoint(set(t_tuple[j])) for (i,j) in it.combinations(range(len(t_tuple)), 2)]):
+                if all([set(t_tuple[i]).isdisjoint(set(t_tuple[j])) for (i,j) in
+                it.combinations(range(len(t_tuple)), 2)]):
                     #Add clause forbidding red K_t